Linux System Administrators’ Guide 0.6

Linux System Administrators’ Guide 0.6

Lars Wirzenius

The Linux Documentation Project

This is version 0.6 of the Linux System Administrators’ Guide. Published November 15, 1997.

The L^{A}TEX source code and other machine readable formats can be found on the Internet via anonymous ftp on sunsite.unc.edu, in the directory /pub/Linux/docs/LDP. Also available are at least Postscript and TEX .DVI formats. The official home page for the book is http://www.iki.fi/liw/linux/sag/. The current version can always be found at that location.

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1993–1997 Lars Wirzenius.

Trademarks are owned by their owners.

Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies.

Permission is granted to process the document source code through TEX or other formatters and print the results, and distribute the printed document, provided the printed document carries copying permission notice identical to this one, including the references to where the source code can be found and the official home page.

Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one.

Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions.

The author would appreciate a notification of modifications, translations, and printed versions. Thank you.

This page is dedicated to a future dedication.

In the mean time… I’d like someone who knows him let Terry Pratchett know that his way of using footnotes is rather inspiring.

Short Contents

1 Introduction 1 2 Overview of a Linux System 6 3 Overview of the Directory Tree 14 4 Using Disks and Other Storage Media 24 5 Memory Management 56 6 Boots And Shutdowns 64

7 init  72

8 Logging In And Out 78 9 Managing user accounts 84 10 Backups 90 11 Keeping Time 100 A Measuring Holes 105 B Glossary (DRAFT) 106

Bibliography 107

ii

Index 108

Contents

1 Introduction 1

2 Overview of a Linux System 6 2.1 Various parts of an operating system . . . . . . . . . . . . . . . . . . 6 2.2 Important parts of the kernel . . . . . . . . . . . . . . . . . . . . . . 7 2.3 Major services in a UNIX system . . . . . . . . . . . . . . . . . . . . 8

2.3.1 init . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.3.2 Logins from terminals . . . . . . . . . . . . . . . . . . . . . . 9 2.3.3 Syslog . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.3.4 Periodic command execution: cron and at . . . . . . . . . . . 10 2.3.5 Graphical user interface . . . . . . . . . . . . . . . . . . . . . 10 2.3.6 Networking . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 2.3.7 Network logins . . . . . . . . . . . . . . . . . . . . . . . . . . 11 2.3.8 Network file systems . . . . . . . . . . . . . . . . . . . . . . . 11 2.3.9 Mail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 2.3.10 Printing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

2.4 The filesystem layout . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

3 Overview of the Directory Tree 14 3.1 Background . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 3.2 The root filesystem . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

3.2.1 The /etc directory . . . . . . . . . . . . . . . . . . . . . . . . 18

3.2.2 The /dev directory . . . . . . . . . . . . . . . . . . . . . . . . 20 3.3 The /usr filesystem . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 3.4 The /var filesystem . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 3.5 The /proc filesystem . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

4 Using Disks and Other Storage Media 24 4.1 Two kinds of devices . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 4.2 Hard disks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 4.3 Floppies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 4.4 CD-ROM’s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 4.5 Tapes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 4.6 Formatting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 4.7 Partitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

4.7.1 The MBR, boot sectors and partition table . . . . . . . . . . . 34 4.7.2 Extended and logical partitions . . . . . . . . . . . . . . . . . 35 4.7.3 Partition types . . . . . . . . . . . . . . . . . . . . . . . . . . 36 4.7.4 Partitioning a hard disk . . . . . . . . . . . . . . . . . . . . . 36 4.7.5 Device files and partitions . . . . . . . . . . . . . . . . . . . . 38

4.8 Filesystems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 4.8.1 What are filesystems? . . . . . . . . . . . . . . . . . . . . . . 38 4.8.2 Filesystems galore . . . . . . . . . . . . . . . . . . . . . . . . . 39 4.8.3 Which filesystem should be used? . . . . . . . . . . . . . . . . 42 4.8.4 Creating a filesystem . . . . . . . . . . . . . . . . . . . . . . . 42 4.8.5 Mounting and unmounting . . . . . . . . . . . . . . . . . . . . 43 4.8.6 Checking filesystem integrity with fsck . . . . . . . . . . . . . 47 4.8.7 Checking for disk errors with badblocks . . . . . . . . . . . . 49 4.8.8 Fighting fragmentation . . . . . . . . . . . . . . . . . . . . . . 49 4.8.9 Other tools for all filesystems . . . . . . . . . . . . . . . . . . 50 4.8.10 Other tools for the ext2 filesystem . . . . . . . . . . . . . . . . 50 4.9 Disks without filesystems . . . . . . . . . . . . . . . . . . . . . . . . . 52 4.10 Allocating disk space . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

4.10.1 Partitioning schemes . . . . . . . . . . . . . . . . . . . . . . . 53 4.10.2 Space requirements . . . . . . . . . . . . . . . . . . . . . . . . 54 4.10.3 Examples of hard disk allocation . . . . . . . . . . . . . . . . 54 4.10.4 Adding more disk space for Linux . . . . . . . . . . . . . . . . 55 4.10.5 Tips for saving disk space . . . . . . . . . . . . . . . . . . . . 55

5 Memory Management 56 5.1 What is virtual memory? . . . . . . . . . . . . . . . . . . . . . . . . . 56 5.2 Creating a swap space . . . . . . . . . . . . . . . . . . . . . . . . . . 57 5.3 Using a swap space . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 5.4 Sharing swap spaces with other operating systems . . . . . . . . . . . 60 5.5 Allocating swap space . . . . . . . . . . . . . . . . . . . . . . . . . . 60 5.6 The buffer cache . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

6 Boots And Shutdowns 64 6.1 An overview of boots and shutdowns . . . . . . . . . . . . . . . . . . 64 6.2 The boot process in closer look . . . . . . . . . . . . . . . . . . . . . 65 6.3 More about shutdowns . . . . . . . . . . . . . . . . . . . . . . . . . . 68 6.4 Rebooting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 6.5 Single user mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 6.6 Emergency boot floppies . . . . . . . . . . . . . . . . . . . . . . . . . 70

7 init  72

7.1 init comes first . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 7.2 Configuring init to start getty: the /etc/inittab file . . . . . . . . 73 7.3 Run levels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 7.4 Special configuration in /etc/inittab . . . . . . . . . . . . . . . . . 75 7.5 Booting in single user mode . . . . . . . . . . . . . . . . . . . . . . . 76 8 Logging In And Out 78

8.1 Logins via terminals . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 8.2 Logins via the network . . . . . . . . . . . . . . . . . . . . . . . . . . 79 8.3 What login does . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 8.4 X and xdm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 8.5 Access control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 8.6 Shell startup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81

9 Managing user accounts 84 9.1 What’s an account? . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 9.2 Creating a user . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85

9.2.1 /etc/passwd and other informative files . . . . . . . . . . . . 85 9.2.2 Picking numeric user and group ids . . . . . . . . . . . . . . . 86 9.2.3 Initial environment: /etc/skel . . . . . . . . . . . . . . . . . 86 9.2.4 Creating a user by hand . . . . . . . . . . . . . . . . . . . . . 87

9.3 Changing user properties . . . . . . . . . . . . . . . . . . . . . . . . . 88 9.4 Removing a user . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 9.5 Disabling a user temporarily . . . . . . . . . . . . . . . . . . . . . . . 89

10 Backups 90 10.1 On the importance of being backed up . . . . . . . . . . . . . . . . . 90 10.2 Selecting the backup medium . . . . . . . . . . . . . . . . . . . . . . 91 10.3 Selecting the backup tool . . . . . . . . . . . . . . . . . . . . . . . . . 92 10.4 Simple backups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

10.4.1 Making backups with tar . . . . . . . . . . . . . . . . . . . . 93 10.4.2 Restoring files with tar . . . . . . . . . . . . . . . . . . . . . 95

10.5 Multilevel backups . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 10.6 What to back up . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 10.7 Compressed backups . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 11 Keeping Time 100

11.1 Time zones . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 11.2 The hardware and software clocks . . . . . . . . . . . . . . . . . . . . 101 11.3 Showing and setting time . . . . . . . . . . . . . . . . . . . . . . . . . 102 11.4 When the clock is wrong . . . . . . . . . . . . . . . . . . . . . . . . . 103

A Measuring Holes 105

B Glossary (DRAFT) 106

Bibliography 107

Index 108

Chapter 1

Introduction

In the beginning, the file was without form, and void; and emptiness was upon the face of the bits. And the Fingers of the Author moved upon the face of the keyboard. And the Author

said, Let there be words, and there were words.

This manual, the Linux System Administrators’ Guide, describes the system administration aspects of using Linux. It is intended for people who know next to nothing about system administration (as in “what is it?”), but who have already mastered at least the basics of normal usage. This manual also doesn’t tell you how to install Linux; that is described in the Installation and Getting Started document. See below for more information about Linux manuals.

System administration is all the things that one has to do to keep a computer system in a useable shape. It includes things like backing up files (and restoring them if necessary), installing new programs, creating accounts for users (and deleting them when no longer needed), making certain that the filesystem is not corrupted, and so on. If a computer were, say, a house, system administration would be called maintenance, and would include cleaning, fixing broken windows, and other such things. System administration is not called maintenance, because that would be too simple.^{1}

The structure of this manual is such that many of the chapters should be usable independently, so that if you need information about, say, backups, you can read just that chapter.^{2} This hopefully makes the book easier to use as a reference manual, and makes it possible to read just a small part when needed, instead of having to read everything. However, this manual is first and foremost a tutorial, and a reference manual only as a lucky coincidence.

1There are some people who do call it that, but that’s just because they have never read this manual, poor things.

2If you happen to be reading a version that has a chapter on backups, that is.

This manual is not intended to be used completely by itself. Plenty of the rest of the Linux documentation is also important for system administrators. After all, a system administrator is just a user with special privileges and duties. A very important resource are the manual pages, which should always be consulted when a command is not familiar.

While this manual is targeted at Linux, a general principle has been that it should be useful with other UNIX based operating systems as well. Unfortunately, since there is so much variance between different versions of UNIX in general, and in system administration in particular, there is little hope to cover all variants. Even covering all possibilities for Linux is difficult, due to the nature of its development.

There is no one official Linux distribution, so different people have different setups, and many people have a setup they have built up themselves. This book is not targeted at any one distribution, even though I use the Debian GNU/Linux system almost exclusively. When possible, I have tried to point out differences, and explain several alternatives.

I have tried to describe how things work, rather than just listing “five easy steps” for each task. This means that there is much information here that is not necessary for everyone, but those parts are marked as such and can be skipped if you use a preconfigured system. Reading everything will, naturally, increase your understanding of the system and should make using and administering it more pleasant.

Like all other Linux related development, the work was done on a volunteer basis: I did it because I thought it might be fun and because I felt it should be done. However, like all volunteer work, there is a limit to how much effort I have been able to spend, and also on how much knowledge and experience I have. This means that the manual is not necessarily as good as it would be if a wizard had been paid handsomely to write it and had spent a few years to perfect it. I think, of course, that it is pretty nice, but be warned.

One particular point where I have cut corners is that I have not covered very thoroughly many things that are already well documented in other freely available manuals. This applies especially to program specific documentation, such as all the details of using mkfs). I only describe the purpose of the program, and as much of its usage as is necessary for the purposes of this manual. For further information, I refer the gentle reader to these other manuals. Usually, all of the referred to documentation is part of the full Linux documentation set.

While I have tried to make this manual as good as possible, I would really like to hear from you if you have any ideas on how to make it better. Bad language, factual errors, ideas for new areas to cover, rewritten sections, information about how various UNIX versions do things, I am interested in all of it. My contact information is

available via the World Wide Web at http://www.iki.fi/liw/mail-to-lasu.html.

You need to read this web page to bypass my junkmail filters.

Many people have helped me with this book, directly or indirectly. I would like to especially thank Matt Welsh for inspiration and LDP leadership, Andy Oram for getting me to work again with much-valued feedback, Olaf Kirch for showing me that it can be done, and Adam Richter at Yggdrasil and others for showing me that other people can find it interesting as well.

Stephen Tweedie, H. Peter Anvin, Rémy Card, Theodore Ts’o, and Stephen Tweedie have let me borrow their work^{3} (and thus make the book look thicker and much more impressive). I am most grateful for this, and very apologetic for the earlier versions that sometimes lacked proper attribution.

In addition, I would like to thank Mark Komarinski for sending his material in 1993 and the many system administration columns in Linux Journal. They are quite informative and inspirational.

Many useful comments have been sent by a large number of people. My miniature black hole of an archive doesn’t let me find all their names, but some of them are, in alphabetical order: Paul Caprioli, Ales Cepek, Marie-France Declerfayt, Dave Dobson, Olaf Flebbe, Helmut Geyer, Larry Greenfield and his father, Stephen Harris, Jyrki Havia, Jim Haynes, York Lam, Timothy Andrew Lister, Jim Lynch, Michael J. Micek, Jacob Navia, Dan Poirier, Daniel Quinlan, Jouni K Seppänen, Philippe Steindl, G.B. Stotte. My apologies to anyone I have forgotten.

Typographical conventions

Bold Used to mark new concepts, WARNINGS, and keywords in a language.

italics Used for emphasis in text, and occasionally for quotes or introductions

3A comparison between the xia and ext2 filesystems, the device list and a description of the ext2 filesystem. These aren’t part of the book any more.

at the beginning of a section.

slanted Used to mark meta-variables in the text, especially in representations of the command line. For example,

               ls -l foo

where foo would “stand for” a filename, such as /bin/cp.

Typewriter Used to represent screen interaction, as in

                    $ ls -l /bin/cp
                    -rwxr-xr-x  1 root    wheel    12104 Sep 25 15:53 /bin/cp

Also used for code examples, whether it is C code, a shell script, or something else, and to display general files, such as configuration files. When necessary for clarity’s sake, these examples or figures will be enclosed in thin boxes.

Key Represents a key to press. You will often see it in this form:

Press return to continue.

The Linux Documentation Project

The Linux Documentation Project, or LDP, is a loose team of writers, proofreaders, and editors who are working together to provide complete documentation for the Linux operating system. The overall coordinator of the project is Greg Hankins.

This manual is one in a set of several being distributed by the LDP, including a Linux Users’ Guide, System Administrators’ Guide, Network Administrators’ Guide, and Kernel Hackers’ Guide. These manuals are all available in L^{A}TEX source format, .dvi format, and postscript output by anonymous FTP from sunsite.unc.edu, in

the directory /pub/Linux/docs/LDP.

We encourage anyone with a penchant for writing or editing to join us in improving Linux documentation. If you have Internet e-mail access, you can contact Greg

Hankins at gregh@sunsite.unc.edu.

CONTENTS 5

The LDP Rhyme^{4}

A wondrous thing, and beautiful, ’tis to write, a book.

I’d like to sing, of the sweat, the blood and tear, which it also took.

It started back in, nineteen-ninety-two, when users whined, "we can nothing do!"

They wanted to know, what their problem was, and how to fix it (by yesterday).

We put the answers in, a Linux f-a-q, hoped to get away, from any more writin’.

"That’s too long, it’s hard to search, and we don’t read it, any-which-way!"

Then a few of us, joined together (virtually, you know), to start the LDP.

4The author wishes to remain anonymous. It was posted to the LDP mailing list by Matt Welsh.

We started to write, or plan, at least, several books, one for every need.

The start was fun,

a lot of talk, an outline, then a slew.

Then silence came, the work began, some wrote less, others more.

A blank screen, oh its horrible, it sits there, laughs in the face.

We still await, the final day, when everything, will be done.

Until then, all we have, is a draft, for you to comment on.

Overview of a Linux System

God looked over everything he had made,

and saw that it was very good. Genesis 1:31

This chapter gives an overview of a Linux system. First, the major services provided by the operating system are described. Then, the programs that implement these services are described with a considerable lack of detail. The purpose of this chapter is to give an understanding of the system as a whole, so that each part is described in detail elsewhere.

2.1 Various parts of an operating system

A UNIX operating system consists of a kernel and some system programs. There are also some application programs for doing work. The kernel is the heart of the operating system^{1}. It keeps track of files on the disk, starts programs and runs them concurrently, assigns memory and other resources to various processes, receives packets from and sends packets to the network, and so on. The kernel does very little by itself, but it provides tools with which all services can be built. It also prevents anyone from accessing the hardware directly, forcing everyone to use the tools it provides. This way the kernel provides some protection for users from each other. The tools provided by the kernel are used via system calls; see manual page section 2 for more information on these.

1In fact, it is often mistakenly considered to be the operating system itself, but it is not. An operating system provides many more services than a plain kernel.

The system programs use the tools provided by the kernel to implement the various services required from an operating system. System programs, and all other programs, run ‘on top of the kernel’, in what is called the user mode. The difference between system and application programs is one of intent: applications are intended for getting useful things done (or for playing, if it happens to be a game), whereas system programs are needed to get the system working. A word processor is an application; telnet is a system program. The difference is often somewhat blurry, however, and is important only to compulsive categorizers.

An operating system can also contain compilers and their corresponding libraries (GCC and the C library in particular under Linux), although not all programming languages need be part of the operating system. Documentation, and sometimes even games, can also be part of it. Traditionally, the operating system has been defined by the contents of the installation tape or disks; with Linux it is not as clear since it is spread all over the FTP sites of the world.

2.2 Important parts of the kernel

The Linux kernel consists of several important parts: process management, memory management, hardware device drivers, filesystem drivers, network management, and various other bits and pieces. Figure 2.1 shows some of them.

Probably the most important parts of the kernel (nothing else works without them) are memory management and process management. Memory management takes care of assigning memory areas and swap space areas to processes, parts of the kernel, and for the buffer cache. Process management creates processes, and implements multitasking by switching the active process on the processor.

At the lowest level, the kernel contains a hardware device driver for each kind of hardware it supports. Since the world is full of different kinds of hardware, the number of hardware device drivers is large. There are often many otherwise similar pieces of hardware that differ in how they are controlled by software. The similarities make it possible to have general classes of drivers that support similar operations; each member of the class has the same interface to the rest of the kernel but differs in what it needs to do to implement them. For example, all disk drivers look alike to the rest of the kernel, i.e., they all have operations like ‘initialize the drive’, ‘read sector N’, and ‘write sector N’.

Some software services provided by the kernel itself have similar properties, and can therefore be abstracted into classes. For example, the various network protocols

Figure 2.1: Some of the more important parts of the Linux kernel.

have been abstracted into one programming interface, the BSD socket library. Another example is the virtual filesystem (VFS) layer that abstracts the filesystem operations away from their implementation. Each filesystem type provides an implementation of each filesystem operation. When some entity tries to use a filesystem, the request goes via the VFS, which routes the request to the proper filesystem driver.

2.3 Major services in a UNIX system

This section describes some of the more important UNIX services, but without much detail. They are described more thoroughly in later chapters.

2.3.1 init

The single most important service in a UNIX system is provided by init. init is started as the first process of every UNIX system, as the last thing the kernel does when it boots. When init starts, it continues the boot process by doing various startup chores (checking and mounting filesystems, starting daemons, etc).

The exact list of things that init does depends on which flavor it is; there are several to choose from. init usually provides the concept of single user mode, in which no one can log in and root uses a shell at the console; the usual mode is called multiuser mode. Some flavors generalize this as run levels; single and multiuser modes are considered to be two run levels, and there can be additional ones as well, for example, to run X on the console.

In normal operation, init makes sure gettys are working (to allow users to log in), and to adopt orphan processes (processes whose parent has died; in UNIX all processes must be in a single tree, so orphans must be adopted).

When the system is shut down, it is init that is in charge of killing all other processes, unmounting all filesystems and stopping the processor, along with anything else it has been configured to do.

2.3.2 Logins from terminals

Logins from terminals (via serial lines) and the console (when not running X) are provided by the getty program. init starts a separate instance of getty for each terminal for which logins are to be allowed. getty reads the username and runs the login program, which reads the password. If the username and password are correct, login runs the shell. When the shell terminates, i.e., the user logs out, or when login terminated because the username and password didn’t match, init notices this and starts a new instance of getty. The kernel has no notion of logins, this is all handled by the system programs.

2.3.3 Syslog

The kernel and many system programs produce error, warning, and other messages. It is often important that these messages can be viewed later, even much later, so they should be written to a file. The program doing this is syslog. It can be configured to sort the messages to different files according to writer or degree of importance. For example, kernel messages are often directed to a separate file from the others, since kernel messages are often more important and need to be read regularly to spot problems.

2.3.4 Periodic command execution: cron and at

Both users and system administrators often need to run commands periodically. For example, the system administrator might want to run a command to clean the directories with temporary files (/tmp and /var/tmp) from old files, to keep the disks from filling up, since not all programs clean up after themselves correctly.

The cron service is set up to do this. Each user has a crontab, where he lists the commands he wants to execute and the times they should be executed. The cron daemon takes care of starting the commands when specified.

The at service is similar to cron, but it is once only: the command is executed at the given time, but it is not repeated.

2.3.5 Graphical user interface

UNIX and Linux don’t incorporate the user interface into the kernel; instead, they let it be implemented by user level programs. This applies for both text mode and graphical environments.

This arrangement makes the system more flexible, but has the disadvantage that it is simple to implement a different user interface for each program, making the system harder to learn.

The graphical environment primarily used with Linux is called the X Window System (X for short). X also does not implement a user interface; it only implements a window system, i.e., tools with which a graphical user interface can be implemented. The three most popular user interface styles implemented over X are Athena, Motif, and Open Look.

2.3.6 Networking

Networking is the act of connecting two or more computers so that they can communicate with each other. The actual methods of connecting and communicating are slightly complicated, but the end result is very useful.

UNIX operating systems have many networking features. Most basic services— filesystems, printing, backups, etc—can be done over the network. This can make system administration easier, since it allows centralized administration, while still reaping in the benefits of microcomputing and distributed computing, such as lower costs and better fault tolerance.

However, this book merely glances at networking; see the Linux Network Administrators’ Guide for more information, including a basic description of how networks operate.

2.3.7 Network logins

Network logins work a little differently than normal logins. There is a separate physical serial line for each terminal via which it is possible to log in. For each person logging in via the network, there is a separate virtual network connection, and there can be any number of these^{2}. It is therefore not possible to run a separate getty for each possible virtual connection. There are also several different ways to log in via a network, telnet and rlogin being the major ones in TCP/IP networks.

Network logins have, instead of a herd of gettys, a single daemon per way of logging in (telnet and rlogin have separate daemons) that listens for all incoming login attempts. When it notices one, it starts a new instance of itself to handle that single attempt; the original instance continues to listen for other attempts. The new instance works similarly to getty.

2.3.8 Network file systems

One of the more useful things that can be done with networking services is sharing files via a network file system. The one usually used is called the Network File System, or NFS, developed by Sun.

With a network file system any file operations done by a program on one machine are sent over the network to another computer. This fools the program to think that all the files on the other computer are actually on the computer the program is running on. This makes information sharing extremely simple, since it requires no modifications to programs.

2Well, at least there can be many. Network bandwidth still being a scarce resource, there is still some practical upper limit to the number of concurrent logins via one network connection.

2.3.9 Mail

Electronic mail is usually the most important method for communicating via computer. An electronic letter is stored in a file using a special format, and special mail programs are used to send and read the letters.

Each user has an incoming mailbox (a file in the special format), where all new mail is stored. When someone sends mail, the mail program locates the receiver’s mailbox and appends the letter to the mailbox file. If the receiver’s mailbox is in another machine, the letter is sent to the other machine, which delivers it to the mailbox as it best sees fit.

The mail system consists of many programs. The delivery of mail to local or remote mailboxes is done by one program (the mail transfer agent or MTA, e.g., sendmail or smail), while the programs users use are many and varied (mail user agent or MUA, e.g., pine or elm). The mailboxes are usually stored in /var/spool/mail.

2.3.10 Printing

Only one person can use a printer at one time, but it is uneconomical not to share printers between users. The printer is therefore managed by software that implements a print queue: all print jobs are put into a queue and whenever the printer is done with one job, the next one is sent to it automatically. This relieves the users from organizing the print queue and fighting over control of the printer.^{3}

The print queue software also spools the printouts on disk, i.e., the text is kept in a file while the job is in the queue. This allows an application program to spit out the print jobs quickly to the print queue software; the application does not have to wait until the job is actually printed to continue. This is really convenient, since it allows one to print out one version, and not have to wait for it to be printed before one can make a completely revised new version.

2.4 The filesystem layout

The filesystem is divided into many parts; usually along the lines of a root filesystem with /bin, /lib, /etc, /dev, and a few others; a /usr filesystem with programs and unchanging data; a /var filesystem with changing data (such as log files); and a /home filesystem for everyone’s personal files. Depending on the hardware configuration and the decisions of the system administrator, the division can be different; it can even be all in one filesystem.

3Instead, they form a new queue at the printer, waiting for their printouts, since no one ever seems to be able to get the queue software to know exactly when anyone’s printout is really finished. This is a great boost to intra-office social relations.

Chapter 3 describes the filesystem layout in some detail; the Linux Filesystem Standard covers it in somewhat more detail.

Overview of the Directory Tree

Two days later, there was Pooh, sitting on his branch, dangling his legs, and there, beside him, were four

pots of honey… (A.A. Milne)

This chapter describes the important parts of a standard Linux directory tree, based on the FSSTND filesystem standard. It outlines the normal way of breaking the directory tree into separate filesystems with different purposes and gives the motivation behind this particular split. Some alternative ways of splitting are also described.

3.1 Background

This chapter is loosely based on the Linux filesystem standard, FSSTND, version 1.2 (see the bibliography, [Qui95]), which attempts to set a standard for how the directory tree in a Linux system is organized. Such a standard has the advantage that it will be easier to write or port software for Linux, and to administer Linux machines, since everything will be in their usual places. There is no authority behind the standard that forces anyone to comply with it, but it has got the support of most, if not all, Linux distributions. It is not a good idea to break with the FSSTND without very compelling reasons. The FSSTND attempts to follow Unix tradition and current trends, making Linux systems familiar to those with experience with other Unix systems, and vice versa.

This chapter is not as detailed as the FSSTND. A system administrator should also read the FSSTND for a complete understanding.

This chapter does not explain all files in detail. The intention is not to describe every file, but to give an overview of the system from a filesystem point of view. Further information on each file is available elsewhere in this manual or the manual pages.

The full directory tree is intended to be breakable into smaller parts, each on its own disk or partition, to accomodate to disk size limits and to ease backup and other system administration. The major parts are the root, /usr, /var, and /home filesystems (see figure 3.1). Each part has a different purpose. The directory tree has been designed so that it works well in a network of Linux machines which may share some parts of the filesystems over a read-only device (e.g., a CD-ROM), or over the network with NFS.

Figure 3.1: Parts of a Unix directory tree. Dashed lines indicate partition limits.

The roles of the different parts of the directory tree are described below.

• The root filesystem is specific for each machine (it is generally stored on a local disk, although it could be a ramdisk or network drive as well) and contains the files that are necessary for booting the system up, and to bring it up to such a state that the other filesystems may be mounted. The contents of the root filesystem will therefore be sufficient for the single user state. It will also contain tools for fixing a broken system, and for recovering lost files from backups.

• The /usr filesystem contains all commands, libraries, manual pages, and other unchanging files needed during normal operation. No files in /usr should be specific for any given machine, nor should they be modified during normal use. This allows the files to be shared over the network, which can be cost-effective since it saves disk space (there can easily be hundreds of megabytes in /usr), and can make administration easier (only the master /usr needs to be changed when updating an application, not each machine separately). Even if the filesystem is on a local disk, it could be mounted read-only, to lessen the chance of filesystem corruption during a crash.

• The /var filesystem contains files that change, such as spool directories (for mail, news, printers, etc), log files, formatted manual pages, and temporary files. Traditionally everything in /var has been somewhere below /usr, but that made it impossible to mount /usr read-only.

• The /home filesystem contains the users’ home directories, i.e., all the real data on the system. Separating home directories to their own directory tree or filesystem makes backups easier; the other parts often do not have to be backed up, or at least not as often (they seldom change). A big /home might have to be broken on several filesystems, which requires adding an extra naming level below /home,

    e.g., /home/students and /home/staff.

Although the different parts have been called filesystems above, there is no requirement that they actually be on separate filesystems. They could easily be kept in a single one if the system is a small single-user system and the user wants to keep things simple. The directory tree might also be divided into filesystems differently, depending on how large the disks are, and how space is allocated for various purposes. The important part, though, is that all the standard names work; even if, say, /var and /usr are actually on the same partition, the names /usr/lib/libc.a and /var/adm/messages must work, for example by moving files below /var into /usr/var, and making /var a symlink to /usr/var.

The Unix filesystem structure groups files according to purpose, i.e., all commands are in one place, all data files in another, documentation in a third, and so on. An alternative would be to group files files according to the program they belong to, i.e., all Emacs files would be in one directory, all TEX in another, and so on. The problem with the latter approach is that it makes it difficult to share files (the program directory often contains both static and shareable and changing and non-shareable files), and sometimes to even find the files (e.g., manual pages in a huge number of places, and making the manual page programs find all of them is a maintenance nightmare).

3.2 The root filesystem

The root filesystem should generally be small, since it contains very critical files and a small, infrequently modified filesystem has a better chance of not getting corrupted. A corrupted root filesystem will generally mean that the system becomes unbootable except with special measures (e.g., from a floppy), so you don’t want to risk it.

The root directory generally doesn’t contain any files, except perhaps the standard boot image for the system, usually called /vmlinuz. All other files are in subdirectories in the root filesystems:

/bin Commands needed during bootup that might be used by normal users (probably after bootup).

/sbin Like /bin, but the commands are not intended for normal users, although they may use them if necessary and allowed.

/etc Configuration files specific to the machine.

/root The home directory for user root.

/lib Shared libraries needed by the programs on the root filesystem.

/lib/modules Loadable kernel modules, especially those that are needed to boot the system when recovering from disasters (e.g., network and filesystem drivers).

/dev Device files.

/tmp Temporary files. Programs running after bootup should use /var/tmp, not /tmp, since the former is probably on a disk with more space.

/boot Files used by the bootstrap loader, e.g., LILO. Kernel images are often kept here instead of in the root directory. If there are many kernel images, the directory can easily grow rather big, and it might be better to keep it in a separate filesystem. Another reason would be to make sure the kernel images are within the first 1024 cylinders of an IDE disk.

/mnt Mount point for temporary mounts by the system administrator. Programs aren’t supposed to mount on /mnt automatically. /mnt might be divided into subdirectories (e.g., /mnt/dosa might be the floppy drive using an MS-DOS filesystem, and /mnt/exta might be the same with an ext2 filesystem).

/proc, /usr, /var, /home Mount points for the other filesystems.

3.2.1 The /etc directory

The /etc directory contains a lot of files. Some of them are described below. For others, you should determine which program they belong to and read the manual page for that program. Many networking configuration files are in /etc as well, and are described in the Networking Administrators’ Guide.

/etc/rc or /etc/rc.d or /etc/rc?.d Scripts or directories of scripts to run at startup or when changing the run level. See the chapter on init for further information.

/etc/passwd The user database, with fields giving the username, real name, home directory, encrypted password, and other information about each user. The format is documented in the passwd manual page.

/etc/fdprm Floppy disk parameter table. Describes what different floppy disk formats look like. Used by setfdprm. See the setfdprm manual page for more information.

/etc/fstab Lists the filesystems mounted automatically at startup by the mount -a command (in /etc/rc or equivalent startup file). Under Linux, also contains information about swap areas used automatically by swapon -a. See section 4.8.5 and the mount manual page for more information.

/etc/group Similar to /etc/passwd, but describes groups instead of users. See the group manual page for more information.

/etc/inittab Configuration file for init.

/etc/issue Output by getty before the login prompt. Usually contains a short description or welcoming message to the system. The contents are up to the system administrator.

/etc/magic The configuration file for file. Contains the descriptions of various file formats based on which file guesses the type of the file. See the magic and file manual pages for more information.

/etc/motd The message of the day, automatically output after a successful login. Contents are up to the system administrator. Often used for getting information to every user, such as warnings about planned downtimes.

/etc/mtab List of currently mounted filesystems. Initially set up by the scripts, and updated automatically by the mount command. Used when a list of mounted filesystems is needed, e.g., by the df command.

/etc/shadow Shadow password file on systems with shadow password software installed. Shadow passwords move the encrypted password from /etc/passwd into /etc/shadow; the latter is not readable by anyone except root. This makes it harder to crack passwords.

/etc/login.defs Configuration file for the login command.

/etc/printcap Like /etc/termcap, but intended for printers. Different syntax.

/etc/profile, /etc/csh.login, /etc/csh.cshrc Files executed at login or startup

time by the Bourne or C shells. These allow the system administrator to set global defaults for all users. See the manual pages for the respective shells.

/etc/securetty Identifies secure terminals, i.e., the terminals from which root is allowed to log in. Typically only the virtual consoles are listed, so that it becomes impossible (or at least harder) to gain superuser privileges by breaking into a system over a modem or a network.

/etc/shells Lists trusted shells. The chsh command allows users to change their login shell only to shells listed in this file. ftpd, the server process that provides FTP services for a machine, will check that the user’s shell is listed in /etc/shells and will not let people log in unles the shell is listed there.

/etc/termcap The terminal capability database. Describes by what “escape sequences” various terminals can be controlled. Programs are written so that instead of directly outputting an escape sequence that only works on a particular brand of terminal, they look up the correct sequence to do whatever it is they want to do in /etc/termcap. As a result most programs work with most kinds of terminals. See the termcap, curs_termcap, and terminfo manual pages for more information.

3.2.2 The /dev directory

The /dev directory contains the special device files for all the devices. The device files are named using special conventions; these are described in the Device list (see [Anv]). The device files are created during installation, and later with the /dev/MAKEDEV script. The /dev/MAKEDEV.local is a script written by the system administrator that creates local-only device files or links (i.e., those that are not part of the standard MAKEDEV, such as device files for some non-standard device driver).

3.3 The /usr filesystem

The /usr filesystem is often large, since all programs are installed there. All files in /usr usually come from a Linux distribution; locally installed programs and other stuff goes below /usr/local. This makes it possible to update the system from a new version of the distribution, or even a completely new distribution, without having to install all programs again. Some of the subdirectories of /usr are listed below (some of the less important directories have been dropped; see the FSSTND for more information).

/usr/X11R6 The X Window System, all files. To simplify the development and installation of X, the X files have not been integrated into the rest of the system. There is a directory tree below /usr/X11R6 similar to that below /usr itself.

/usr/X386 Similar to /usr/X11R6, but for X11 Release 5.

/usr/bin Almost all user commands. Some commands are in /bin or in /usr/local/bin.

/usr/sbin System administration commands that are not needed on the root filesystem, e.g., most server programs.

/usr/man, /usr/info, /usr/doc Manual pages, GNU Info documents, and miscellaneous other documentation files, respectively.

/usr/include Header files for the C programming language. This should actually be below /usr/lib for consistency, but the tradition is overwhelmingly in support for this name.

/usr/lib Unchanging data files for programs and subsystems, including some site-wide configuration files. The name lib comes from library; origi-

nally libraries of programming subroutines were stored in /usr/lib.

/usr/local The place for locally installed software and other files.

3.4 The /var filesystem

The /var contains data that is changed when the system is running normally. It is specific for each system, i.e., not shared over the network with other computers.

/var/catman A cache for man pages that are formatted on demand. The source for manual pages is usually stored in /usr/man/man*; some manual pages might come with a pre-formatted version, which is stored in /usr/man/cat*. Other manual pages need to be formatted when they are first viewed; the formatted version is then stored in /var/man so that the next person to view the same page won’t have to wait for it to be formatted. (/var/catman is often cleaned in the same way temporary directories are cleaned.)

/var/lib Files that change while the system is running normally.

/var/local Variable data for programs that are installed in /usr/local (i.e., programs that have been installed by the system administrator). Note that even locally installed programs should use the other /var directories if they are appropriate, e.g., /var/lock.

/var/lock Lock files. Many programs follow a convention to create a lock file in /var/lock to indicate that they are using a particular device or file. Other programs will notice the lock file and won’t attempt to use the device or file.

/var/log Log files from various programs, especially login (/var/log/wtmp, which logs all logins and logouts into the system) and syslog (/var/log/messages, where all kernel and system program message are usually stored). Files in /var/log can often grow indefinitely, and may require cleaning at regular intervals.

/var/run Files that contain information about the system that is valid until the system is next booted. For example, /var/run/utmp contains information about people currently logged in.

/var/spool Directories for mail, news, printer queues, and other queued work. Each different spool has its own subdirectory below /var/spool, e.g., the mailboxes of the users are in /var/spool/mail.

/var/tmp Temporary files that are large or that need to exist for a longer time than what is allowed for /tmp. (Although the system administrator might not allow very old files in /var/tmp either.)

3.5 The /proc filesystem

The /proc filesystem contains a illusionary filesystem. It does not exist on a disk. Instead, the kernel creates it in memory. It is used to provide information about the system (originally about processes, hence the name). Some of the more important files and directories are explained below. The /proc filesystem is described in more detail in the proc manual page.

/proc/1 A directory with information about process number 1. Each process has a directory below /proc with the name being its process identification number.

/proc/cpuinfo Information about the processor, such as its type, make, model, and perfomance.

/proc/devices List of device drivers configured into the currently running kernel.

/proc/dma Shows which DMA channels are being used at the moment.

/proc/filesystems Filesystems configured into the kernel.

/proc/interrupts Shows which interrupts are in use, and how many of each there have been.

/proc/ioports Which I/O ports are in use at the moment.

/proc/kcore An image of the physical memory of the system. This is exactly the same size as your physical memory, but does not really take up that much memory; it is generated on the fly as programs access it. (Remember: unless you copy it elsewhere, nothing under /proc takes up any disk space at all.)

/proc/kmsg Messages output by the kernel. These are also routed to syslog. /proc/ksyms Symbol table for the kernel. /proc/loadavg The ‘load average’ of the system; three meaningless indicators of

how much work the system has to do at the moment.

/proc/meminfo Information about memory usage, both physical and swap. /proc/modules Which kernel modules are loaded at the moment. /proc/net Status information about network protocols. /proc/self A symbolic link to the process directory of the program that is looking

at /proc. When two processes look at /proc, they get different links. This is mainly a convenience to make it easier for programs to get at their process directory.

/proc/stat Various statistics about the system, such as the number of page faults

since the system was booted.

/proc/uptime The time the system has been up. /proc/version The kernel version. Note that while the above files tend to be easily readable text files, they can sometimes be formatted in a way that is not easily digestable. There are many commands that do little more than read the above files and format them for easier understanding. For example, the free program reads /proc/meminfo and converts the amounts given in bytes to kilobytes (and adds a little more information, as well).

Using Disks and Other Storage Media

On a clear disk you can seek forever.

When you install or upgrade your system, you need to do a fair amount of work on your disks. You have to make filesystems on your disks so that files can be stored on them and reserve space for the different parts of your system.

This chapter explains all these initial activities. Usually, once you get your system set up, you won’t have to go through the work again, except for using floppies. You’ll need to come back to this chapter if you add a new disk or want to fine-tune your disk usage.

The basic tasks in administering disks are:

• Format your disk. This does various things to prepare it for use, such as checking for bad sectors. (Formatting is nowadays not necessary for most hard disks.)

• Partition a hard disk, if you want to use it for several activities that aren’t supposed to interfere with one another. One reason for partitioning is to store different operating systems on the same disk. Another reason is to keep user files separate from system files, which simplifies back-ups and helps protect the system files from corruption.

• Make a filesystem (of a suitable type) on each disk or partition. The disk means nothing to Linux until you make a filesystem; then files can be created and accessed on it.

• Mount different filesystems to form a single tree structure, either automatically, or manually as needed. (Manually mounted filesystems usually need to be unmounted manually as well.)

Chapter 5 contains information about virtual memory and disk caching, of which you also need to be aware when using disks.

This chapter explains what you need to know for hard disks, floppies, CD-ROM’s, and tape drives.

4.1 Two kinds of devices

UNIX, and therefore Linux, recognizes two different kinds of device: random-access block devices (such as disks), and character devices (such as tapes and serial lines), some of which may be serial, and some random-access. Each supported device is represented in the filesystem as a device file. When you read or write a device file, the data comes from or goes to the device it represents. This way no special programs (and no special application programming methodology, such as catching interrupts or polling a serial port) are necessary to access devices; for example, to send a file to the printer, one could just say

$ cat filename > /dev/lp1

     $

and the contents of the file are printed (the file must, of course, be in a form that the printer understands). However, since it is not a good idea to have several people cat their files to the printer at the same time, one usually uses a special program to send the files to be printed (usually lpr). This program makes sure that only one file is being printed at a time, and will automatically send files to the printer as soon as it finishes with the previous file. Something similar is needed for most devices. In fact, one seldom needs to worry about device files at all.

Since devices show up as files in the filesystem (in the /dev directory), it is easy to see just what device files exist, using ls or another suitable command. In the output of ls -l, the first column contains the type of the file and its permissions. For example, inspecting a serial device gives on my system

$ ls -l /dev/cua0

     crw-rw-rw-   1 root     uucp       5,  64 Nov 30  1993 /dev/cua0
     $

The first character in the first column, i.e., ‘c’ in crw-rw-rw- above, tells an informed user the type of the file, in this case a character device. For ordinary files, the first character is ‘-’, for directories it is ‘d’, and for block devices ‘b’; see the ls man page for further information.

Note that usually all device files exist even though the device itself might be not be installed. So just because you have a file /dev/sda, it doesn’t mean that you really do have an SCSI hard disk. Having all the device files makes the installation programs simpler, and makes it easier to add new hardware (there is no need to find out the correct parameters for and create the device files for the new device).

4.2 Hard disks

This subsection introduces terminology related to hard disks. If you already know the terms and concepts, you can skip this subsection.

See figure 4.1 for a schematic picture of the important parts in a hard disk. A hard disk consists of one or more circular platters,^{1} of which either or both surfaces are coated with a magnetic substance used for recording the data. For each surface, there is a read-write head that examines or alters the recorded data. The platters rotate on a common axis; a typical rotation speed is 3600 rotations per minute, although high-performance hard disks have higher speeds. The heads move along the radius of the platters; this movement combined with the rotation of the platters allows the head to access all parts of the surfaces.

The processor (CPU) and the actual disk communicate through a disk controller. This relieves the rest of the computer from knowing how to use the drive, since the controllers for different types of disks can be made to use the same interface towards the rest of the computer. Therefore, the computer can say just “hey disk, gimme what I want”, instead of a long and complex series of electric signals to move the head to the proper location and waiting for the correct position to come under the head and doing all the other unpleasant stuff necessary. (In reality, the interface to the controller is still complex, but much less so than it would otherwise be.) The controller can also do some other stuff, such as caching, or automatic bad sector replacement.

The above is usually all one needs to understand about the hardware. There is also a bunch of other stuff, such as the motor that rotates the platters and moves the heads, and the electronics that control the operation of the mechanical parts, but that is mostly not relevant for understanding the working principle of a hard disk.

1The platters are made of a hard substance, e.g., aluminium, which gives the hard disk its name.

The surfaces are usually divided into concentric rings, called tracks, and these in turn are divided into sectors. This division is used to specify locations on the hard disk and to allocate disk space to files. To find a given place on the hard disk, one might say “surface 3, track 5, sector 7”. Usually the number of sectors is the same for all tracks, but some hard disks put more sectors in outer tracks (all sectors are of the same physical size, so more of them fit in the longer outer tracks). Typically, a sector will hold 512 bytes of data. The disk itself can’t handle smaller amounts of data than one sector.

Figure 4.1: A schematic picture of a hard disk.

Each surface is divided into tracks (and sectors) in the same way. This means that when the head for one surface is on a track, the heads for the other surfaces are also on the corresponding tracks. All the corresponding tracks taken together are called a cylinder. It takes time to move the heads from one track (cylinder) to another, so by placing the data that is often accessed together (say, a file) so that it is within one cylinder, it is not necessary to move the heads to read all of it. This improves performance. It is not always possible to place files like this; files that are stored in several places on the disk are called fragmented.

The number of surfaces (or heads, which is the same thing), cylinders, and sectors vary a lot; the specification of the number of each is called the geometry of a hard disk. The geometry is usually stored in a special, battery-powered memory location called the CMOS RAM, from where the operating system can fetch it during bootup or driver initialization.

Unfortunately, the BIOS^{2} has a design limitation, which makes it impossible to specify a track number that is larger than 1024 in the CMOS RAM, which is too little for a large hard disk. To overcome this, the hard disk controller lies about the geometry, and translates the addresses given by the computer into something that fits reality. For example, a hard disk might have 8 heads, 2048 tracks, and 35 sectors per track^{3}. Its controller could lie to the computer and claim that it has 16 heads, 1024 tracks, and 35 sectors per track, thus not exceeding the limit on tracks, and translates the address that the computer gives it by halving the head number, and doubling the track number. The math can be more complicated in reality, because the numbers are not as nice as here (but again, the details are not relevant for understanding the principle). This translation distorts the operating system’s view of how the disk is organized, thus making it impractical to use the all-data-on-one-cylinder trick to boost performance.

The translation is only a problem for IDE disks. SCSI disks use a sequential sector number (i.e., the controller translates a sequential sector number to a head, cylinder, and sector triplet), and a completely different method for the CPU to talk with the controller, so they are insulated from the problem. Note, however, that the computer might not know the real geometry of an SCSI disk either.

Since Linux often will not know the real geometry of a disk, its filesystems don’t even try to keep files within a single cylinder. Instead, it tries to assign sequentially numbered sectors to files, which almost always gives similar performance. The issue is further complicated by on-controller caches, and automatic prefetches done by the controller.

Each hard disk is represented by a separate device file. There can (usually) be only two or four IDE hard disks. These are known as /dev/hda, /dev/hdb, /dev/hdc, and /dev/hdd, respectively. SCSI hard disks are known as /dev/sda, /dev/sdb, and so on. Similar naming conventions exist for other hard disk types; see [Anv] for more information. Note that the device files for the hard disks give access to the entire disk, with no regard to partitions (which will be discussed below), and it’s easy to mess up the partitions or the data in them if you aren’t careful. The disks’ device files are usually used only to get access to the master boot record (which will also be discussed below).

2The BIOS is some built-in software stored on ROM chips. It takes care, among other things, of the initial stages of booting.

3The numbers are completely imaginary.

4.3 Floppies

A floppy disk consists of a flexible membrane covered on one or both sides with similar magnetic substance as a hard disk. The floppy disk itself doesn’t have a read-write head, that is included in the drive. A floppy corresponds to one platter in a hard disk, but is removable and one drive can be used to access different floppies, whereas the hard disk is one indivisible unit.

Like a hard disk, a floppy is divided into tracks and sectors (and the two corresponding tracks on either side of a floppy form a cylinder), but there are many fewer of them than on a hard disk.

A floppy drive can usually use several different types of disks; for example, a 3\frac{1}{2} inch drive can use both 720 kB and 1.44 MB disks. Since the drive has to operate a bit differently and the operating system must know how big the disk is, there are many device files for floppy drives, one per combination of drive and disk type. Therefore, /dev/fd0H1440 is the first floppy drive (fd0), which must be a 3\frac{1}{2} inch drive, using a 3\frac{1}{2} inch, high density disk (H) of size 1440 kB (1440), i.e., a normal 3\frac{1}{2} inch HD floppy. For more information on the naming conventions for the floppy devices, see [Anv].

The names for floppy drives are complex, however, and Linux therefore has a special floppy device type that automatically detects the type of the disk in the drive. It works by trying to read the first sector of a newly inserted floppy using different floppy types until it finds the correct one. This naturally requires that the floppy is formatted first. The automatic devices are called /dev/fd0, /dev/fd1, and so on.

The parameters the automatic device uses to access a disk can also be set using the program setfdprm. This can be useful if you need to use disks that do not follow any usual floppy sizes, e.g., if they have an unusual number of sectors, or if the autodetecting for some reason fails and the proper device file is missing.

Linux can handle many nonstandard floppy disk formats in addition to all the standard ones. Some of these require using special formatting programs. We’ll skip these disk types for now, but in the mean time you can examine the /etc/fdprm file. It specifies the settings that setfdprm recognizes.

The operating system must know when a disk has been changed in a floppy drive, for example, in order to avoid using cached data from the previous disk. Unfortunately, the signal line that is used for this is sometimes broken, and worse, this won’t always be noticeable when using the drive from within MS-DOS. If you are experiencing weird problems using floppies, this might be the reason. The only way to correct it is to repair the floppy drive.

4.4 CD-ROM’s

A CD-ROM drive uses an optically read, plastic coated disk. The information is recorded on the surface of the disk^{4} in small ‘holes’ aligned along a spiral from the center to the edge. The drive directs a laser beam along the spiral to read the disk. When the laser hits a hole, the laser is reflected in one way; when it hits smooth surface, it is reflected in another way. This makes it easy to code bits, and therefore information. The rest is easy, mere mechanics.

CD-ROM drives are slow compared to hard disks. Whereas a typical hard disk will have an average seek time less than 15 milliseconds, a fast CD-ROM drive can use tenths of a second for seeks. The actual data transfer rate is fairly high at hundreds of kilobytes per second. The slowness means that CD-ROM drives are not as pleasant to use instead of hard disks (some Linux distributions provide ‘live’ filesystems on CD-ROM’s, making it unnecessary to copy the files to the hard disk, making installation easier and saving a lot of hard disk space), although it is still possible. For installing new software, CD-ROM’s are very good, since it maximum speed is not essential during installation.

There are several ways to arrange data on a CD-ROM. The most popular one is specified by the international standard ISO 9660. This standard specifies a very minimal filesystem, which is even more crude than the one MS-DOS uses. On the other hand, it is so minimal that every operating system should be able to map it to its native system.

For normal UNIX use, the ISO 9660 filesystem is not usable, so an extension to the standard has been developed, called the Rock Ridge extension. Rock Ridge allows longer filenames, symbolic links, and a lot of other goodies, making a CD-ROM look more or less like any contemporary UNIX filesystem. Even better, a Rock Ridge filesystem is still a valid ISO 9660 filesystem, making it usable by non-UNIX systems as well. Linux supports both ISO 9660 and the Rock Ridge extensions; the extensions are recognized and used automatically.

The filesystem is only half the battle, however. Most CD-ROM’s contain data that requires a special program to access, and most of these programs do not run under Linux (except, possibly, under dosemu, the Linux MS-DOS emulator).

4That is, the surface inside the disk, on the metal disk inside the plastic coating.

A CD-ROM drive is accessed via the corresponding device file. There are several ways to connect a CD-ROM drive to the computer: via SCSI, via a sound card, or via EIDE. The hardware hacking needed to do this is outside the scope of this book, but the type of connection decides the device file. See [Anv] for enlightment.

4.5 Tapes

A tape drive uses a tape, similar^{5} to cassettes used for music. A tape is serial in nature, which means that in order to get to any given part of it, you first have to go through all the parts in between. A disk can be accessed randomly, i.e., you can jump directly to any place on the disk. The serial access of tapes makes them slow.

On the other hand, tapes are relatively cheap to make, since they do not need to be fast. They can also easily be made quite long, and can therefore contain a large amount of data. This makes tapes very suitable for things like archiving and backups, which do not require large speeds, but benefit from low costs and large storage capacities.

4.6 Formatting

Formatting is the process of writing marks on the magnetic media that are used to mark tracks and sectors. Before a disk is formatted, its magnetic surface is a complete mess of magnetic signals. When it is formatted, some order is brought into the chaos by essentially drawing lines where the tracks go, and where they are divided into sectors. The actual details are not quite exactly like this, but that is irrelevant. What is important is that a disk cannot be used unless it has been formatted.

The terminology is a bit confusing here: in MS-DOS, the word formatting is used to cover also the process of creating a filesystem (which will be discussed below). There, the two processes are often combined, especially for floppies. When the distinction needs to be made, the real formatting is called low-level formatting, while making the filesystem is called high-level formatting. In UNIX circles, the two are called formatting and making a filesystem, so that’s what is used in this book as well.

For IDE and some SCSI disks the formatting is actually done at the factory and doesn’t need to be repeated; hence most people rarely need to worry about it. In fact, formatting a hard disk can cause it to work less well, for example because a disk might need to be formatted in some very special way to allow automatic bad sector replacement to work.

5But completely different, of course.

Disks that need to be or can be formatted often require a special program anyway, because the interface to the formatting logic inside the drive is different from drive to drive. The formatting program is often either on the controller BIOS, or is supplied as an MS-DOS program; neither of these can easily be used from within Linux.

During formatting one might encounter bad spots on the disk, called bad blocks or bad sectors. These are sometimes handled by the drive itself, but even then, if more of them develop, something needs to be done to avoid using those parts of the disk. The logic to do this is built into the filesystem; how to add the information into the filesystem is described below. Alternatively, one might create a small partition that covers just the bad part of the disk; this approach might be a good idea if the bad spot is very large, since filesystems can sometimes have trouble with very large bad areas.

Floppies are formatted with fdformat. The floppy device file to use is given as the parameter. For example, the following command would format a high density, 3\frac{1}{2} inch floppy in the first floppy drive:

$ fdformat /dev/fd0H1440

     Double-sided, 80 tracks, 18 sec/track. Total capacity 1440 kB.
     Formatting ... done
     Verifying ... done
     $

Note that if you want to use an autodetecting device (e.g., /dev/fd0), you must set the parameters of the device with setfdprm first. To achieve the same effect as above, one would have to do the following:

$ setfdprm /dev/fd0 1440/1440

$ fdformat /dev/fd0

     Double-sided, 80 tracks, 18 sec/track. Total capacity 1440 kB.
     Formatting ... done
     Verifying ... done
     $

It is usually more convenient to choose the correct device file that matches the type of the floppy. Note that it is unwise to format floppies to contain more information than what they are designed for.

fdformat will also validate the floppy, i.e., check it for bad blocks. It will try a bad block several times (you can usually hear this, the drive noise changes dramatically). If the floppy is only marginally bad (due to dirt on the read/write head, some errors are false signals), fdformat won’t complain, but a real error will abort the validation process. The kernel will print log messages for each I/O error it finds; these will go to the console or, if syslog is being used, to the file /usr/adm/messages. fdformat itself won’t tell where the error is (one usually doesn’t care, floppies are cheap enough that a bad one is automatically thrown away).

$ fdformat /dev/fd0H1440

     Double-sided, 80 tracks, 18 sec/track. Total capacity 1440 kB.
     Formatting ... done
     Verifying ... read: Unknown error
     $

The badblocks command can be used to search any disk or partition for bad blocks (including a floppy). It does not format the disk, so it can be used to check even existing filesystems. The example below checks a 3\frac{1}{2} inch floppy with two bad blocks.

$ badblocks /dev/fd0H1440 1440

     718
     719
     $

badblocks outputs the block numbers of the bad blocks it finds. Most filesystems can avoid such bad blocks. They maintain a list of known bad blocks, which is initialized when the filesystem is made, and can be modified later. The initial search for bad blocks can be done by the mkfs command (which initializes the filesystem), but later checks should be done with badblocks and the new blocks should be added with fsck. We’ll describe mkfs and fsck later.

Many modern disks automatically notice bad blocks, and attempt to fix them by using a special, reserved good block instead. This is invisible to the operating system. This feature should be documented in the disk’s manual, if you’re curious if it is happening. Even such disks can fail, if the number of bad blocks grows too large, although chances are that by then the disk will be so rotten as to be unusable.

4.7 Partitions

A hard disk can be divided into several partitions. Each partition functions as if it were a separate hard disk. The idea is that if you have one hard disk, and want to have, say, two operating systems on it, you can divide the disk into two partitions. Each operating system uses its partition as it wishes and doesn’t touch the other one’s. This way the two operating systems can co-exist peacefully on the same hard disk. Without partitions one would have to buy a hard disk for each operating system.

Floppies are not partitioned. There is no technical reason against this, but since they’re so small, partitions would be useful only very rarely. CD-ROM’s are usually also not partitioned, since it’s easier to use them as one big disk, and there is seldom a need to have several operating systems on one.

4.7.1 The MBR, boot sectors and partition table

The information about how a hard disk has been partitioned is stored in its first sector (that is, the first sector of the first track on the first disk surface). The first sector is the master boot record (MBR) of the disk; this is the sector that the BIOS reads in and starts when the machine is first booted. The master boot record contains a small program that reads the partition table, checks which partition is active (that is, marked bootable), and reads the first sector of that partition, the partition’s boot sector (the MBR is also a boot sector, but it has a special status and therefore a special name). This boot sector contains another small program that reads the first part of the operating system stored on that partition (assuming it is bootable), and then starts it.

The partitioning scheme is not built into the hardware, or even into the BIOS. It is only a convention that many operating systems follow. Not all operating systems do follow it, but they are the exceptions. Some operating systems support partitions, but they occupy one partition on the hard disk, and use their internal partitioning method within that partition. The latter type exists peacefully with other operating systems (including Linux), and does not require any special measures, but an operating system that doesn’t support partitions cannot co-exist on the same disk with any other operating system.

As a safety precaution, it is a good idea to write down the partition table on a piece of paper, so that if it ever corrupts you don’t have to lose all your files. (A bad partition table can be fixed with fdisk). The relevant information is given by the fdisk -l command:

     $ fdisk -l /dev/hda
     Disk /dev/hda: 15 heads, 57 sectors, 790 cylinders
     Units = cylinders of 855 * 512 bytes
        Device Boot  Begin   Start     End  Blocks   Id  System
     /dev/hda1           1       1      24   10231+  82  Linux swap
     /dev/hda2          25      25      48   10260   83  Linux native
     /dev/hda3          49      49     408  153900   83  Linux native
     /dev/hda4         409     409     790  163305    5  Extended
     /dev/hda5         409     409     744  143611+  83  Linux native
     /dev/hda6         745     745     790   19636+  83  Linux native
     $

4.7.2 Extended and logical partitions

The original partitioning scheme for PC hard disks allowed only four partitions. This quickly turned out to be too little in real life, partly because some people want more than four operating systems (Linux, MS-DOS, OS/2, Minix, FreeBSD, NetBSD, or Windows/NT, to name a few), but primarily because sometimes it is a good idea to have several partitions for one operating system. For example, swap space is usually best put in its own partition for Linux instead of in the main Linux partition for reasons of speed (see below).

To overcome this design problem, extended partitions were invented. This trick allows partitioning a primary partition into sub-partitions. The primary partition thus subdivided is the extended partition; the subpartitions are logical partitions. They behave like primary^{6} partitions, but are created differently. There is no speed difference between them.

The partition structure of a hard disk might look like that in figure 4.2. The disk is divided into three primary partitions, the second of which is divided into two logical partitions. Part of the disk is not partitioned at all. The disk as a whole and each primary partition has a boot sector.

6Illogical?

Figure 4.2: A sample hard disk partitioning.

4.7.3 Partition types

The partition tables (the one in the MBR, and the ones for extended partitions) contain one byte per partition that identifies the type of that partition. This attempts to identify the operating system that uses the partition, or what it uses it for. The purpose is to make it possible to avoid having two operating systems accidentally using the same partition. However, in reality, operating systems do not really care about the partition type byte; e.g., Linux doesn’t care at all what it is. Worse, some of them use it incorrectly; e.g., at least some versions of DR-DOS ignore the most significant bit of the byte, while others don’t.

There is no standardization agency to specify what each byte value means, but some commonly accepted ones are included in in table 4.1. The same list is available in the Linux fdisk program.

4.7.4 Partitioning a hard disk

There are many programs for creating and removing partitions. Most operating systems have their own, and it can be a good idea to use each operating system’s own, just in case it does something unusual that the others can’t. Many of the programs are called fdisk, including the Linux one, or variations thereof. Details on using the Linux fdisk are given on its man page. The cfdisk command is similar to fdisk, but has a nicer (full screen) user interface.

When using IDE disks, the boot partition (the partition with the bootable kernel

Table 4.1: Partition types (from the Linux fdisk program).

0 Empty  40 Venix 80286  94 Amoeba BBT
1 DOS 12-bit FAT  51 Novell?  a5 BSD/386
2 XENIX root  52 Microport  b7 BSDI fs
3 XENIX usr  63 GNU HURD b8 BSDI swap
4 DOS 16-bit <32M 64 Novell  c7 Syrinx
5 Extended 75 PC/IX db CP/M
6 DOS 16-bit ≥32M 80 Old MINIX  e1 DOS access
7 OS/2 HPFS  81 Linux/MINIX e3 DOS R/O
8 AIX  82 Linux swap  f2 DOS secondary
9 AIX bootable  83 Linux native  ff BBT
a OS/2 Boot Manag 93 Amoeba

image files) must be completely within the first 1024 cylinders. This is because the disk is used via the BIOS during boot (before the system goes into protected mode), and BIOS can’t handle more than 1024 cylinders. It is sometimes possible to use a boot partition that is only partly within the first 1024 cylinders. This works as long as all the files that are read with the BIOS are within the first 1024 cylinders. Since this is difficult to arrange, it is a very bad idea to do it; you never know when a kernel update or disk defragmentation will result in an unbootable system. Therefore, make sure your boot partition is completely within the first 1024 cylinders.

Some newer versions of the BIOS and IDE disks can, in fact, handle disks with more than 1024 cylinders. If you have such a system, you can forget about the problem; if you aren’t quite sure of it, put it within the first 1024 cylinders.

Each partition should have an even number of sectors, since the Linux filesystems use a 1 kB block size, i.e., two sectors. An odd number of sectors will result in the last sector being unused. This won’t result in any problems, but it is ugly, and some versions of fdisk will warn about it.

Changing a partition’s size usually requires first backing up everything you want to save from that partition (preferably the whole disk, just in case), deleting the partition, creating new partition, then restoring everything to the new partition. If the partition is growing, you may need to adjust the sizes (and backup and restore) of the adjoining partitions as well.

Since changing partition sizes is painful, it is preferable to get the partitions right the first time, or have an effective and easy to use backup system. If you’re installing from a media that does not require much human intervention (say, from CD-ROM, as opposed to floppies), it is often easy to play with different configuration at first. Since you don’t already have data to back up, it is not so painful to modify partition sizes several times.

There is a program for MS-DOS, called fips, which resizes an MS-DOS partition without requiring the backup and restore, but for other filesystems it is still necessary.

4.7.5 Device files and partitions

Each partition and extended partition has its own device file. The naming convention for these files is that a partition’s number is appended after the name of the whole disk, with the convention that 1–4 are primary partitions (regardless of how many primary partitions there are) and 5–8 are logical partitions (regardless of within which primary partition they reside). For example, /dev/hda1 is the first primary partition on the first IDE hard disk, and /dev/sdb7 is the third extended partition on the second SCSI hard disk. The device list in [Anv] gives more information.

4.8 Filesystems

4.8.1 What are filesystems?

A filesystem is the methods and data structures that an operating system uses to keep track of files on a disk or partition; that is, the way the files are organized on the disk. The word is also used to refer to a partition or disk that is used to store the files or the type of the filesystem. Thus, one might say “I have two filesystems” meaning one has two partitions on which one stores files, or that one is using the “extended filesystem”, meaning the type of the filesystem.

The difference between a disk or partition and the filesystem it contains is important. A few programs (including, reasonably enough, programs that create filesystems) operate directly on the raw sectors of a disk or partition; if there is an existing file system there it will be destroyed or seriously corrupted. Most programs operate on a filesystem, and therefore won’t work on a partition that doesn’t contain one (or that contains one of the wrong type).

Before a partition or disk can be used as a filesystem, it needs to be initialized, and the bookkeeping data structures need to be written to the disk. This process is called making a filesystem.

Most UNIX filesystem types have a similar general structure, although the exact details vary quite a bit. The central concepts are superblock, inode, data block, directory block, and indirection block. The superblock contains information about the filesystem as a whole, such as its size (the exact information here depends on the filesystem). An inode contains all information about a file, except its name. The name is stored in the directory, together with the number of the inode. A directory entry consists of a filename and the number of the inode which represents the file. The inode contains the numbers of several data blocks, which are used to store the data in the file. There is space only for a few data block numbers in the inode, however, and if more are needed, more space for pointers to the data blocks is allocated dynamically. These dynamically allocated blocks are indirect blocks; the name indicates that in order to find the data block, one has to find its number in the indirect block first.

UNIX filesystems usually allow one to create a hole in a file (this is done with lseek; check the manual page), which means that the filesystem just pretends that at a particular place in the file there is just zero bytes, but no actual disk sectors are reserved for that place in the file (this means that the file will use a bit less disk space). This happens especially often for small binaries, Linux shared libraries, some databases, and a few other special cases. (Holes are implemented by storing a special value as the address of the data block in the indirect block or inode. This special address means that no data block is allocated for that part of the file, ergo, there is a hole in the file.)

Holes are moderately useful. On the author’s system, a simple measurement showed a potential for about 4 MB of savings through holes of about 200 MB total used disk space. That system, however, contains relatively few programs and no database files. The measurement tool is described in appendix A.

4.8.2 Filesystems galore

Linux supports several types of filesystems. As of this writing the most important ones are:

minix The oldest, presumed to be the most reliable, but quite limited in features (some time stamps are missing, at most 30 character filenames) and restricted in capabilities (at most 64 MB per filesystem).

xia A modified version of the minix filesystem that lifts the limits on the filenames and filesystem sizes, but does not otherwise introduce new features. It is not very popular, but is reported to work very well. ext2 The most featureful of the native Linux filesystems, currently also the

most popular one. It is designed to be easily upwards compatible, so that new versions of the filesystem code do not require re-making the existing filesystems.

ext An older version of ext2 that wasn’t upwards compatible. It is hardly ever used in new installations any more, and most people have converted

           to ext2.

In addition, support for several foreign filesystem exists, to make it easier to exchange files with other operating systems. These foreign filesystems work just like native ones, except that they may be lacking in some usual UNIX features, or have curious limitations, or other oddities.

msdos Compatibility with MS-DOS (and OS/2 and Windows NT) FAT filesystems.

umsdos Extends the msdos filesystem driver under Linux to get long filenames, owners, permissions, links, and device files. This allows a normal msdos filesystem to be used as if it were a Linux one, thus removing the need for a separate partition for Linux.

iso9660 The standard CD-ROM filesystem; the popular Rock Ridge extension to the CD-ROM standard that allows longer file names is supported automatically.

nfs A networked filesystem that allows sharing a filesystem between many computers to allow easy access to the files from all of them.

hpfs The OS/2 filesystem.

sysv SystemV/386, Coherent, and Xenix filesystems.

The choice of filesystem to use depends on the situation. If compatibility or other reasons make one of the non-native filesystems necessary, then that one must be used. If one can choose freely, then it is probably wisest to use ext2, since it has all the features but does not suffer from lack of performance.

There is also the proc filesystem, usually accessible as the /proc directory, which is not really a filesystem at all, even though it looks like one. The proc filesystem makes it easy to access certain kernel data structures, such as the process list (hence

4.8. Filesystems 41

the name). It makes these data structures look like a filesystem, and that filesystem can be manipulated with all the usual file tools. For example, to get a listing of all processes one might use the command

$ ls -l /proc

total 0
dr-xr-xr-x   4 root     root            0 Jan 31 20:37 1
dr-xr-xr-x   4 liw      users           0 Jan 31 20:37 63
dr-xr-xr-x   4 liw      users           0 Jan 31 20:37 94
dr-xr-xr-x   4 liw      users           0 Jan 31 20:37 95
dr-xr-xr-x   4 root     users           0 Jan 31 20:37 98
dr-xr-xr-x   4 liw      users           0 Jan 31 20:37 99
-r--r--r--   1 root     root            0 Jan 31 20:37 devices
-r--r--r--   1 root     root            0 Jan 31 20:37 dma
-r--r--r--   1 root     root            0 Jan 31 20:37 filesystems
-r--r--r--   1 root     root            0 Jan 31 20:37 interrupts
-r--------   1 root     root      8654848 Jan 31 20:37 kcore
-r--r--r--   1 root     root            0 Jan 31 11:50 kmsg
-r--r--r--   1 root     root            0 Jan 31 20:37 ksyms
-r--r--r--   1 root     root            0 Jan 31 11:51 loadavg
-r--r--r--   1 root     root            0 Jan 31 20:37 meminfo
-r--r--r--   1 root     root            0 Jan 31 20:37 modules
dr-xr-xr-x   2 root     root            0 Jan 31 20:37 net
dr-xr-xr-x   4 root     root            0 Jan 31 20:37 self
-r--r--r--   1 root     root            0 Jan 31 20:37 stat
-r--r--r--   1 root     root            0 Jan 31 20:37 uptime
-r--r--r--   1 root     root            0 Jan 31 20:37 version
$

(There will be a few extra files that don’t correspond to processes, though. The above example has been shortened.)

Note that even though it is called a filesystem, no part of the proc filesystem touches any disk. It exists only in the kernel’s imagination. Whenever anyone tries to look at any part of the proc filesystem, the kernel makes it look as if the part existed somewhere, even though it doesn’t. So, even though there is a multi-megabyte /proc/kcore file, it doesn’t take any disk space.

4.8.3 Which filesystem should be used?

There is usually little point in using many different filesystems. Currently, ext2fs is the most popular one, and it is probably the wisest choice. Depending on the overhead for bookkeeping structures, speed, (perceived) reliability, compatibility, and various other reasons, it may be advisable to use another file system. This needs to be decided on a case-by-case basis.

4.8.4 Creating a filesystem

Filesystems are created, i.e., initialized, with the mkfs command. There is actually a separate program for each filesystem type. mkfs is just a front end that runs the appropriate program depending on the desired filesystem type. The type is selected with the -t fstype option.

The programs called by mkfs have slightly different command line interfaces. The common and most important options are summarized below; see the manual pages for more.

-t fstype Select the type of the filesystem.

-c Search for bad blocks and initialize the bad block list accordingly. -l filename Read the initial bad block list from the file filename.

To create an ext2 filesystem on a floppy, one would give the following commands:

$ fdformat -n /dev/fd0H1440

     Double-sided, 80 tracks, 18 sec/track. Total capacity 1440 kB.
     Formatting ... done

$ badblocks /dev/fd0H1440 1440 > bad-blocks $ mkfs -t ext2 -l bad-blocks /dev/fd0H1440

     mke2fs 0.5a, 5-Apr-94 for EXT2 FS 0.5, 94/03/10
     360 inodes, 1440 blocks
     72 blocks (5.00%) reserved for the super user
     First data block=1
     Block size=1024 (log=0)
     Fragment size=1024 (log=0)
     1 block group
     8192 blocks per group, 8192 fragments per group
     360 inodes per group

     Writing inode tables: done
     Writing superblocks and filesystem accounting information: done
     $

First, the floppy was formatted (the -n option prevents validation, i.e., bad block checking). Then bad blocks were searched with badblocks, with the output redirected to a file, bad-blocks. Finally, the filesystem was created, with the bad block list initialized by whatever badblocks found.

The -c option could have been used with mkfs instead of badblocks and a separate file. The example below does that.

$ mkfs -t ext2 -c /dev/fd0H1440

     mke2fs 0.5a, 5-Apr-94 for EXT2 FS 0.5, 94/03/10
     360 inodes, 1440 blocks
     72 blocks (5.00%) reserved for the super user
     First data block=1
     Block size=1024 (log=0)
     Fragment size=1024 (log=0)
     1 block group
     8192 blocks per group, 8192 fragments per group
     360 inodes per group
     Checking for bad blocks (read-only test): done
     Writing inode tables: done
     Writing superblocks and filesystem accounting information: done
     $

The -c is more convenient than a separate use of badblocks, but badblocks is necessary for checking after the filesystem has been created.

The process to prepare filesystems on hard disks or partitions is the same as for floppies, except that the formatting isn’t needed.

4.8.5 Mounting and unmounting

Before one can use a filesystem, it has to be mounted. The operating system then does various bookkeeping things to make sure that everything works. Since all files in UNIX are in a single directory tree, the mount operation will make it look like the contents of the new filesystem are the contents of an existing subdirectory in some already mounted filesystem.

For example, figure 4.3 shows three separate filesystems, each with their own root directory. When the last two filesystems are mounted below /home and /usr, respectively, on the first filesystem, we can get a single directory tree, as in figure 4.4.

Figure 4.3: Three separate filesystems.

Figure 4.4: /home and /usr have been mounted.

The mounts could be done as in the following example:

$ mount /dev/hda2 /home $ mount /dev/hda3 /usr

     $

The mount command takes two arguments. The first one is the device file corresponding to the disk or partition containing the filesystem. The second one is the directory below which it will be mounted. After these commands the contents of the two filesystems look just like the contents of the /home and /usr directories, respectively. One would then say that “/dev/hda2 is mounted on /home”, and similarly for /usr. To look at either filesystem, one would look at the contents of the directory on which it has been mounted, just as if it were any other directory. Note the difference between the device file, /dev/hda2, and the mounted-on directory, /home. The device file gives access to the raw contents of the disk, the mounted-on directory gives access to the files on the disk. The mounted-on directory is called the mount point.

Linux supports many filesystem types. mount tries to guess the type of the filesystem. You can also use the -t fstype option to specify the type directly; this is sometimes necessary, since the heuristics mount uses do not always work. For example, to mount an MS-DOS floppy, you could use the following command:

$ mount -t msdos /dev/fd0 /floppy

     $

The mounted-on directory need not be empty, although it must exist. Any files in it, however, will be inaccessible by name while the filesystem is mounted. (Any files that have already been opened will still be accessible. Files that have hard links from other directories can be accessed using those names.) There is no harm done with this, and it can even be useful. For instance, some people like to have /tmp and /var/tmp synonymous, and make /tmp be a symbolic link to /var/tmp. When the system is booted, before the /usr filesystem is mounted, a /var/tmp directory residing on the root filesystem is used instead. When /usr is mounted, it will make the /var/tmp directory on the root filesystem inaccessible. If /var/tmp didn’t exist on the root filesystem, it would be impossible to use temporary files before mounting

/var.

If you don’t intend to write anything to the filesystem, use the -r switch for mount to do a readonly mount. This will make the kernel stop any attempts at writing to the filesystem, and will also stop the kernel from updating file access times in the inodes. Read-only mounts are necessary for unwritable media, e.g., CD-ROM’s.

The alert reader has already noticed a slight logistical problem. How is the first filesystem (called the root filesystem, because it contains the root directory) mounted, since it obviously can’t be mounted on another filesystem? Well, the answer is that it is done by magic.^{7} The root filesystem is magically mounted at boot time, and one can rely on it to always be mounted—if the root filesystem can’t be mounted, the system does not boot. The name of the filesystem that is magically mounted as root is either compiled into the kernel, or set using LILO or rdev.

The root filesystem is usually first mounted readonly. The startup scripts will then run fsck to verify its validity, and if there are no problems, they will re-mount it so that writes will also be allowed. fsck must not be run on a mounted filesystem, since any changes to the filesystem while fsck is running will cause trouble. Since the root filesystem is mounted readonly while it is being checked, fsck can fix any problems without worry, since the remount operation will flush any metadata that the filesystem keeps in memory.

7For more information, see the kernel source or the Kernel Hackers’ Guide.

On many systems there are other filesystems that should also be mounted automatically at boot time. These are specified in the /etc/fstab file; see the fstab man page for details on the format. The details of exactly when the extra filesystems are mounted depend on many factors, and can be configured by each administrator if need be. When the chapter on booting is finished, you may read all about it there.

When a filesystem no longer needs to be mounted, it can be unmounted with umount^{8}. umount takes one argument: either the device file or the mount point. For example, to unmount the directories of the previous example, one could use the commands

$ umount /dev/hda2

$ umount /usr

     $

See the man page for further instructions on how to use the command. It is imperative that you always unmount a mounted floppy. Don’t just pop the floppy out of the drive! Because of disk caching, the data is not necessarily written to the floppy until you unmount it, so removing the floppy from the drive too early might cause the contents to become garbled. If you only read from the floppy, this is not very likely, but if you write, even accidentally, the result may be catastrophic.

Mounting and unmounting requires super user privileges, i.e., only root can do it. The reason for this is that if any user can mount a floppy on any directory, then it is rather easy to create a floppy with, say, a Trojan horse disguised as /bin/sh, or any other often used program. However, it is often necessary to allow users to use floppies, and there are several ways to do this:

• Give the users the root password. This is obviously bad security, but is the easiest solution. It works well if there is no need for security anyway, which is the case on many non-networked, personal systems.

• Use a program such as sudo to allow users to use mount. This is still bad security, but doesn’t directly give super user privileges to everyone.^{9}

• Make the users use mtools, a package for manipulating MS-DOS filesystems, without mounting them. This works well if MS-DOS floppies are all that is needed, but is rather awkward otherwise.

8It should of course be unmount, but the n mysteriously disappeared in the 70’s, and hasn’t been seen since. Please return it to Bell Labs, NJ, if you find it.

9It requires several seconds of hard thinking on the users’ behalf.

• List the floppy devices and their allowable mount points together with the suitable options in /etc/fstab.

The last alternative can be implemented by adding a line like the following to the

/etc/fstab file:
     /dev/fd0 /floppy msdos user,noauto 0 0

The columns are: device file to mount, directory to mount on, filesystem type, options, backup frequency (used by dump), and fsck pass number (to specify the order in which filesystems should be checked upon boot; 0 means no check).

The noauto option stops this mount to be done automatically when the system is started (i.e., it stops mount -a from mounting it). The user option allows any user to mount the filesystem, and, because of security reasons, disallows execution of programs (normal or setuid) and interpretation of device files from the mounted filesystem. After this, any user can mount a floppy with an msdos filesystem with the following command:

$ mount /floppy

     $

The floppy can (and needs to, of course) be unmounted with the corresponding umount command.

If you want to provide access to several types of floppies, you need to give several mount points. The settings can be different for each mount point. For example, to give access to both MS-DOS and ext2 floppies, you could have the following to lines

in /etc/fstab:
     /dev/fd0    /dosfloppy    msdos   user,noauto  0  0
     /dev/fd0    /ext2floppy   ext2    user,noauto  0  0

For MS-DOS filesystems (not just floppies), you probably want to restrict access to it by using the uid, gid, and umask filesystem options, described in detail on the mount manual page. If you aren’t careful, mounting an MS-DOS filesystem gives everyone at least read access to the files in it, which is not a good idea.

4.8.6 Checking filesystem integrity with fsck

Filesystems are complex creatures, and as such, they tend to be somewhat error-prone. A filesystem’s correctness and validity can be checked using the fsck command. It can be instructed to repair any minor problems it finds, and to alert the user if there any unrepairable problems. Fortunately, the code to implement filesystems is debugged quite effectively, so there are seldom any problems at all, and they are usually caused by power failures, failing hardware, or operator errors; for example, by not shutting down the system properly.

Most systems are setup to run fsck automatically at boot time, so that any errors are detected (and hopefully corrected) before the system is used. Use of a corrupted filesystem tends to make things worse: if the data structures are messed up, using the filesystem will probably mess them up even more, resulting in more data loss. However, fsck can take a while to run on big filesystems, and since errors almost never occur if the system has been shut down properly, a couple of tricks are used to avoid doing the checks in such cases. The first is that if the file /etc/fastboot exists, no checks are made. The second is that the ext2 filesystem has a special marker in its superblock that tells whether the filesystem was unmounted properly after the previous mount. This allows e2fsck (the version of fsck for the ext2 filesystem) to avoid checking the filesystem if the flag indicates that the unmount was done (the assumption being that a proper unmount indicates no problems). Whether the /etc/fastboot trick works on your system depends on your startup scripts, but the ext2 trick works every time you use e2fsck—it has to be explicitly bypassed with an option to e2fsck to be avoided. (See the e2fsck man page for details on how.)

The automatic checking only works for the filesystems that are mounted automatically at boot time. Use fsck manually to check other filesystems, e.g., floppies.

If fsck finds unrepairable problems, you need either in-depth knowlege of how filesystems work in general, and the type of the corrupt filesystem in particular, or good backups. The latter is easy (although sometimes tedious) to arrange, the former can sometimes be arranged via a friend, the Linux newsgroups and mailing lists, or some other source of support, if you don’t have the know-how yourself. I’d like to tell you more about it, but my lack of education and experience in this regard hinders me. The debugfs program by Theodore T’so should be useful.

fsck must only be run on unmounted filesystems, never on mounted filesystems (with the exception of the read-only root during startup). This is because it accesses the raw disk, and can therefore modify the filesystem without the operating system realizing it. There will be trouble, if the operating system is confused.

4.8.7 Checking for disk errors with badblocks

It can be a good idea to periodically check for bad blocks. This is done with the badblocks command. It outputs a list of the numbers of all bad blocks it can find. This list can be fed to fsck to be recorded in the filesystem data structures so that the operating system won’t try to use the bad blocks for storing data. The following example will show how this could be done.

$ badblocks /dev/fd0H1440 1440 > bad-blocks $ fsck -t ext2 -l bad-blocks /dev/fd0H1440

     Parallelizing fsck version 0.5a (5-Apr-94)
     e2fsck 0.5a, 5-Apr-94 for EXT2 FS 0.5, 94/03/10
     Pass 1: Checking inodes, blocks, and sizes
     Pass 2: Checking directory structure
     Pass 3: Checking directory connectivity
     Pass 4: Check reference counts.
     Pass 5: Checking group summary information.
     /dev/fd0H1440: ***** FILE SYSTEM WAS MODIFIED *****
     /dev/fd0H1440: 11/360 files, 63/1440 blocks
     $

If badblocks reports a block that was already used, e2fsck will try to move the block to another place. If the block was really bad, not just marginal, the contents of the file may be corrupted.

4.8.8 Fighting fragmentation

When a file is written to disk, it can’t always be written in consecutive blocks. A file that is not stored in consecutive blocks is fragmented. It takes longer to read a fragmented file, since the disk’s read-write head will have to move more. It is desireable to avoid fragmentation, although it is less of a problem in a system with a good buffer cache with read-ahead.

The ext2 filesystem attempts to keep fragmentation at a minimum, by keeping all blocks in a file close together, even if they can’t be stored in consecutive sectors. Ext2 effectively always allocates the free block that is nearest to other blocks in a file. For ext2, it is therefore seldom necessary to worry about fragmentation. There is a program for defragmenting an ext2 filesystem, see [TV] in the bibliography.

There are many MS-DOS defragmentation programs that move blocks around in the filesystem to remove fragmentation. For other filesystems, defragmentation must be done by backing up the filesystem, re-creating it, and restoring the files from backups. Backing up a filesystem before defragmening is a good idea for all filesystems, since many things can go wrong during the defragmentation.

4.8.9 Other tools for all filesystems

Some other tools are also useful for managing filesystems. df shows the free disk space on one or more filesystems; du shows how much disk space a directory and all its files contain. These can be used to hunt down disk space wasters.

sync forces all unwritten blocks in the buffer cache (see section 5.6) to be written to disk. It is seldom necessary to do this by hand; the daemon process update does this automatically. It can be useful in catastrophies, for example if update or its helper process bdflush dies, or if you must turn off power now and can’t wait for

update to run.

4.8.10 Other tools for the ext2 filesystem

In addition to the filesystem creator (mke2fs) and checker (e2fsck) accessible directly or via the filesystem type independent front ends, the ext2 filesystem has some additional tools that can be useful.

tune2fs adjusts filesystem parameters. Some of the more interesting parameters are:

• A maximal mount count. e2fsck enforces a check when filesystem has been mounted too many times, even if the clean flag is set. For a system that is used for developing or testing the system, it might be a good idea to reduce this limit.

• A maximal time between checks. e2fsck can also enforce a maximal time between two checks, even if the clean flag is set, and the filesystem hasn’t been mounted very often. This can be disabled, however.

• Number of blocks reserved for root. Ext2 reserves some blocks for root so that if the filesystem fills up, it is still possible to do system administration without having to delete anything. The reserved amount is by default 5%, which on most disks isn’t enough to be wasteful. However, for floppies there is no point in reserving any blocks.

dumpe2fs 0.5b, 11-Mar-95 for EXT2 FS 0.5a, 94/10/23
Filesystem magic number:  0xEF53
Filesystem state:         clean
Errors behavior:          Continue
Inode count:              360
Block count:              1440
Reserved block count:     72
Free blocks:              1133
Free inodes:              326
First block:              1
Block size:               1024
Fragment size:            1024
Blocks per group:         8192
Fragments per group:      8192
Inodes per group:         360
Last mount time:          Tue Aug  8 01:52:52 1995
Last write time:          Tue Aug  8 01:53:28 1995
Mount count:              3
Maximum mount count:      20
Last checked:             Tue Aug  8 01:06:31 1995
Check interval:           0
Reserved blocks uid:      0 (user root)
Reserved blocks gid:      0 (group root)
Group 0:
  Block bitmap at 3, Inode bitmap at 4, Inode table at 5
  1133 free blocks, 326 free inodes, 2 directories
  Free blocks: 307-1439
  Free inodes: 35-360

Figure 4.5: Sample output from dumpe2fs

See the tune2fs manual page for more information.

dumpe2fs shows information about an ext2 filesystem, mostly from the superblock. Figure 4.5 shows a sample output. Some of the information in the output is technical and requires understanding of how the filesystem works (see appendix ??), but much of it is readily understandable even for layadmins.

debugfs is a filesystem debugger. It allows direct access to the filesystem data structures stored on disk and can thus be used to repair a disk that is so broken that fsck can’t fix it automatically. It has also been known to be used to recover deleted files. However, debugfs very much requires that you understand what you’re doing; a failure to understand can destroy all your data.

dump and restore can be used to back up an ext2 filesystem. They are ext2 specific versions of the traditional UNIX backup tools. See chapter 10 for more information on backups.

4.9 Disks without filesystems

Not all disks or partitions are used as filesystems. A swap partition, for example, will not have a filesystem on it. Many floppies are used in a tape-drive emulating fashion, so that a tar or other file is written directly on the raw disk, without a filesystem. Linux boot floppies don’t contain a filesystem, only the raw kernel.

Avoiding a filesystem has the advantage of making more of the disk usable, since a filesystem always has some bookkeeping overhead. It also makes the disks more easily compatible with other systems: for example, the tar file format is the same on all systems, while filesystems are different on most systems. You will quickly get used to disks without filesystems if you need them. Bootable Linux floppies also do not necessarily have a filesystem, although that is also possible.

One reason to use raw disks is to make image copies of them. For instance, if the disk contains a partially damaged filesystem, it is a good idea to make an exact copy of it before trying to fix it, since then you can start again if your fixing breaks things even more. One way to do this is to use dd:

$ dd if=/dev/fd0H1440 of=floppy-image

     2880+0 records in
     2880+0 records out

$ dd if=floppy-image of=/dev/fd0H1440

     2880+0 records in
     2880+0 records out
     $

The first dd makes an exact image of the floppy to the file floppy-image, the second one writes the image to the floppy. (The user has presumably switched the floppy before the second command. Otherwise the command pair is of doubtful usefulness.)

4.10 Allocating disk space

4.10.1 Partitioning schemes

It is not easy to partition a disk in the best possible way. Worse, there is no universally correct way to do it; there are too many factors involved.

The traditional way is to have a (relatively) small root filesystem, which contains /bin, /etc, /dev, /lib, /tmp, and other stuff that is needed to get the system up and running. This way, the root filesystem (in its own partition or on its own disk) is all that is needed to bring up the system. The reasoning is that if the root filesystem is small and is not heavily used, it is less likely to become corrupt when the system crashes, and you will therefore find it easier to fix any problems caused by the crash. Then you create separate partitions or use separate disks for the directory tree below /usr, the users’ home directories (often under /home), and the swap space. Separating the home directories (with the users’ files) in their own partition makes backups easier, since it is usually not necessary to backup programs (which reside below /usr). In a networked environment it is also possible to share /usr among several machines (e.g., by using NFS), thereby reducing the total disk space required by several tens or hundreds of megabytes times the number of machines.

The problem with having many partitions is that it splits the total amount of free disk space into many small pieces. Nowadays, when disks and (hopefully) operating systems are more reliable, many people prefer to have just one partition that holds all their files. On the other hand, it can be less painful to back up (and restore) a small partition.

For a small hard disk (assuming you don’t do kernel development), the best way to go is probably to have just one partition. For large hard disks, it is probably better to have a few large partitions, just in case something does go wrong. (Note that ‘small’ and ‘large’ are used in a relative sense here; your needs for disk space decide what the threshold is.)

If you have several disks, you might wish to have the root filesystem (including /usr) on one, and the users’ home directories on another.

It is a good idea to be prepared to experiment a bit with different partitioning schemes (over time, not just while first installing the system). This is a bit of work, since it essentially requires you to install the system from scratch several times, but it is the only way to be sure you do it right.

4.10.2 Space requirements

The Linux distribution you install will give some indication of how much disk space you need for various configurations. Programs installed separately may also do the same. This will help you plan your disk space usage, but you should prepare for the future and reserve some extra space for things you will notice later that you need.

The amount you need for user files depends on what your users wish to do. Most people seem to need as much space for their files as possible, but the amount they will live happily with varies a lot. Some people do only light text processing and will survive nicely with a few megabytes, others do heavy image processing and will need gigabytes.

By the way, when comparing file sizes given in kilobytes or megabytes and disk space given in megabytes, it can be important to know that the two units can be different. Some disk manufacturers like to pretend that a kilobyte is 1000 bytes and a megabyte is 1000 kilobytes, while all the rest of the computing world uses 1024 for both factors. Therefore, my 345 MB hard disk is really a 330 MB hard disk.^{10}

Swap space allocation is discusses in section 5.5.

4.10.3 Examples of hard disk allocation

I used to have a 109 MB hard disk. Now I am using a 330 MB hard disk. I’ll explain how and why I partitioned these disks.

The 109 MB disk I partitioned in a lot of ways, when my needs and the operating systems I used changed; I’ll explain two typical scenarios. First, I used to run MS-DOS together with Linux. For that, I needed about 20 MB of hard disk, or just enough to have MS-DOS, a C compiler, an editor, a few other utilities, the program I was working on, and enough free disk space to not feel claustrophobic. For Linux, I had a 10 MB swap partition, and the rest, or 79 MB, was a single partition with all the files I had under Linux. I experimented with having separate root, /usr, and /home partitions, but there was never enough free disk space in one piece to do much interesting.

When I didn’t need MS-DOS anymore, I repartitioned the disk so that I had a 12 MB swap partition, and again had the rest as a single filesystem.

The 330 MB disk is partitioned into several partitions, like this:

10Sic transit discus mundi.

5 MB root filesystem
10 MB swap partition
180 MB /usr filesystem
120 MB /home filesystem
15 MB scratch partition

The scratch partition is for playing around with things that require their own partition, e.g., trying different Linux distributions, or comparing speeds of filesystems. When not needed for anything else, it is used as swap space (I like to have a lot of open windows).

4.10.4 Adding more disk space for Linux

Adding more disk space for Linux is easy, at least after the hardware has been properly installed (the hardware installation is outside the scope of this book). You format it if necessary, then create the partitions and filesystem as described above, and add the proper lines to /etc/fstab so that it is mounted automatically.

4.10.5 Tips for saving disk space

The best tip for saving disk space is to avoid installing unnecessary programs. Most Linux distributions have an option to install only part of the packages they contain, and by analyzing your needs you might notice that you don’t need most of them. This will help save a lot of disk space, since many programs are quite large. Even if you do need a particular package or program, you might not need all of it. For example, some on-line documentation might be unnecessary, as might some of the Elisp files for GNU Emacs, some of the fonts for X11, or some of the libraries for programming.

If you cannot uninstall packages, you might look into compression. Compression programs such as gzip or zip will compress (and uncompress) individual files or groups of files. The gzexe system will compress and uncompress programs invisibly to the user (unused programs are compressed, then uncompressed as they are used). The experimental DouBle system will compress all files in a filesystem, invisibly to the programs that use them. (If you are familiar with products such as Stacker for MS-DOS, the principle is the same.)

Memory Management

Minnet, jag har tappat mitt minne,

är jag svensk eller finne kommer inte ihåg… (Bosse Österberg)

This section describes the Linux memory management features, i.e., virtual memory and the disk buffer cache. The purpose and workings and the things the system administrator needs to take into consideration are described.

5.1 What is virtual memory?

Linux supports virtual memory, that is, using a disk as an extension of RAM so that the effective size of usable memory grows correspondingly. The kernel will write the contents of a currently unused block of memory to the hard disk so that the memory can be used for another purpose. When the original contents are needed again, they are read back into memory. This is all made completely transparent to the user; programs running under Linux only see the larger amount of memory available and don’t notice that parts of them reside on the disk from time to time. Of course, reading and writing the hard disk is slower (on the order of a thousand times slower) than using real memory, so the programs don’t run as fast. The part of the hard disk that is used as virtual memory is called the swap space.

Linux can use either a normal file in the filesystem or a separate partition for swap space. A swap partition is faster, but it is easier to change the size of a swap file (there’s no need to repartition the whole hard disk, and possibly install everything from scratch). When you know how much swap space you need, you should go for a swap partition, but if you are uncertain, you can use a swap file first, use the system for a while so that you can get a feel for how much swap you need, and then make a swap partition when you’re confident about its size.

You should also know that Linux allows one to use several swap partitions and/or swap files at the same time. This means that if you only occasionally need an unusual amount of swap space, you can set up an extra swap file at such times, instead of keeping the whole amount allocated all the time.

A note on operating system terminology: computer science usually distinguishes between swapping (writing the whole process out to swap space) and paging (writing only fixed size parts, usually a few kilobytes, at a time). Paging is usually more efficient, and that’s what Linux does, but traditional Linux terminology talks about swapping anyway.^{1}

5.2 Creating a swap space

A swap file is an ordinary file; it is in no way special to the kernel. The only thing that matters to the kernel is that it has no holes, and that it is prepared for use with mkswap. It must reside on a local disk, however; it can’t reside in a filesystem that has been mounted over NFS due to implementation reasons.

The bit about holes is important. The swap file reserves the disk space so that the kernel can quickly swap out a page without having to go through all the things that are necessary when allocating a disk sector to a file. The kernel merely uses any sectors that have already been allocated to the file. Because a hole in a file means that there are no disk sectors allocated (for that place in the file), it is not good for the kernel to try to use them.

One good way to create the swap file without holes is through the following command:

$ dd if=/dev/zero of=/extra-swap bs=1024 count=1024

     1024+0 records in
     1024+0 records out
     $

1Thus quite needlessly annoying a number of computer scientists something horrible.

where /extra-swap is the name of the swap file and the size of is given after the count=. It is best for the size to be a multiple of 4, because the kernel writes out memory pages, which are 4 kilobytes in size. If the size is not a multiple of 4, the last couple of kilobytes may be unused.

A swap partition is also not special in any way. You create it just like any other partition; the only difference is that it is used as a raw partition, that is, it will not contain any filesystem at all. It is a good idea to mark swap partitions as type 82 (Linux swap); this will the make partition listings clearer, even though it is not strictly necessary to the kernel.

After you have created a swap file or a swap partition, you need to write a signature to its beginning; this contains some administrative information and is used by the kernel. The command to do this is mkswap, used like this:

$ mkswap /extra-swap 1024

     Setting up swapspace, size = 1044480 bytes
     $

Note that the swap space is still not in use yet: it exists, but the kernel does not use it to provide virtual memory.

You should be very careful when using mkswap, since it does not check that the file or partition isn’t used for anything else. You can easily overwrite important files and partitions with mkswap! Fortunately, you should only need to use mkswap when you install your system.

The Linux memory manager limits the size of each swap space to about 127 MB (for various technical reasons, the actual limit is (4096−10)×8×4096 = 133890048 bytes, or 127.6875 megabytes). You can, however, use up to 16 swap spaces simultaneously, for a total of almost 2 GB.^{2}

5.3 Using a swap space

An initialized swap space is taken into use with swapon. This command tells the kernel that the swap space can be used. The path to the swap space is given as the argument, so to start swapping on a temporary swap file one might use the following command.

2A gigabyte here, a gigabyte there, pretty soon we start talking about real memory.

$ swapon /extra-swap

     $

Swap spaces can be used automatically by listing them in the /etc/fstab file.

     /dev/hda8        none        swap        sw     0     0
     /swapfile        none        swap        sw     0     0

The startup scripts will run the command swapon -a, which will start swapping on all the swap spaces listed in /etc/fstab. Therefore, the swapon command is usually used only when extra swap is needed.

You can monitor the use of swap spaces with free. It will tell the total amount of swap space used.

$ free

                  total       used       free     shared    buffers
     Mem:         15152      14896        256      12404       2528
     -/+ buffers:            12368       2784
     Swap:        32452       6684      25768
     $

The first line of output (Mem:) shows the physical memory. The total column does not show the physical memory used by the kernel, which is usually about a megabyte. The used column shows the amount of memory used (the second line does not count buffers). The free column shows completely unused memory. The shared column shows the amount of memory shared by several processes; the more, the merrier. The buffers column shows the current size of the disk buffer cache.

That last line (Swap:) shows similar information for the swap spaces. If this line is all zeroes, your swap space is not activated.

The same information is available via top, or using the proc filesystem in file /proc/meminfo. It is currently difficult to get information on the use of a specific swap space.

A swap space can be removed from use with swapoff. It is usually not necessary to do it, except for temporary swap spaces. Any pages in use in the swap space are swapped in first; if there is not sufficient physical memory to hold them, they will then be swapped out (to some other swap space). If there is not enough virtual memory to hold all of the pages Linux will start to thrash; after a long while it should recover, but meanwhile the system is unusable. You should check (e.g., with free) that there is enough free memory before removing a swap space from use.

All the swap spaces that are used automatically with swapon -a can be removed from use with swapoff -a; it looks at the file /etc/fstab to find what to remove. Any manually used swap spaces will remain in use.

Sometimes a lot of swap space can be in use even though there is a lot of free physical memory. This can happen for instance if at one point there is need to swap, but later a big process that occupied much of the physical memory terminates and frees the memory. The swapped-out data is not automatically swapped in until it is needed, so the physical memory may remain free for a long time. There is no need to worry about this, but it can be comforting to know what is happening.

5.4 Sharing swap spaces with other operating systems

Virtual memory is built into many operating systems. Since they each need it only when they are running, i.e., never at the same time, the swap spaces of all but the currently running one are being wasted. It would be more efficient for them to share a single swap space. This is possible, but can require a bit of hacking. The Tips-HOWTO contains some advice on how to implement this.

5.5 Allocating swap space

Some people will tell you that you should allocate twice as much swap space as you have physical memory, but this is a bogus rule. Here’s how to do it properly:

1. Estimate your total memory needs. This is the largest amount of memory you’ll probably need at a time, that is the sum of the memory requirements of all the programs you want to run at the same time. This can be done by running at the same time all the programs you are likely to ever be running at the same time. For instance, if you want to run X, you should allocate about 8 MB for it, gcc wants several megabytes (some files need an unusually large amount, up to tens of megabytes, but usually about four should do), and so on. The kernel will use about a megabyte by itself, and the usual shells and other small utilities perhaps a few hundred kilobytes (say a megabyte together). There is no need to try to be exact, rough estimates are fine, but you might want to be on the pessimistic side.

Remember that if there are going to be several people using the system at the same time, they are all going to consume memory. However, if two people run the same program at the same time, the total memory consumption is usually not double, since code pages and shared libraries exist only once. The free and ps commands are useful for estimating the memory needs.

2. Add some security to the estimate in step 1. This is because estimates of program sizes will probably be wrong, because you’ll probably forget some programs you want to run, and to make certain that you have some extra space just in case. A couple of megabytes should be fine. (It is better to allocate too much than too little swap space, but there’s no need to over-do it and allocate the whole disk, since unused swap space is wasted space; see later about adding more swap.) Also, since it is nicer to deal with even numbers, you can round the value up to the next full megabyte.

3. Based on the computations above, you know how much memory you’ll be needing in total. So, in order to allocate swap space, you just need to subtract the size of your physical memory from the total memory needed, and you know how much swap space you need. (On some versions of UNIX, you need to allocate space for an image of the physical memory as well, so the amount computed in step 2 is what you need and you shouldn’t do the subtraction.)

4. If your calculated swap space is very much larger than your physical memory (more than a couple times larger), you should probably invest in more physical memory, otherwise performance will be too low.

It’s a good idea to have at least some swap space, even if your calculations indicate that you need none. Linux uses swap space somewhat aggressively, so that as much physical memory as possible can be kept free. Linux will swap out memory pages that have not been used, even if the memory is not yet needed for anything. This avoids waiting for swapping when it is needed—the swapping can be done earlier, when the disk is otherwise idle.

Swap space can be divided among several disks. This can sometimes improve performance, depending on the relative speeds of the disks and the access patterns of the disks. You might want to experiment with a few schemes, but be aware that doing the experiments properly is quite difficult. You should not believe claims that any one scheme is superior to any other, since it won’t always be true.

5.6 The buffer cache

Reading from a disk^{3} is very slow compared to accessing (real) memory. In addition, it is common to read the same part of a disk several times during relatively short periods of time. For example, one might first read an e-mail message, then read the letter into an editor when replying to it, then make the mail program read it again when copying it to a folder. Or, consider how often the command ls might be run on a system with many users. By reading the information from disk only once and then keeping it in memory until no longer needed, one can speed up all but the first read. This is called disk buffering, and the memory used for the purpose is called the buffer cache.

Since memory is, unfortunately, a finite, nay, scarce resource, the buffer cache usually cannot be big enough (it can’t hold all the data one ever wants to use). When the cache fills up, the data that has been unused for the longest time is discarded and the memory thus freed is used for the new data.

Disk buffering works for writes as well. On the one hand, data that is written is often soon read again (e.g., a source code file is saved to a file, then read by the compiler), so putting data that is written in the cache is a good idea. On the other hand, by only putting the data into the cache, not writing it to disk at once, the program that writes runs quicker. The writes can then be done in the background, without slowing down the other programs.

Most operating systems have buffer caches (although they might be called something else), but not all of them work according to the above principles. Some are write-through: the data is written to disk at once (it is kept in the cache as well, of course). The cache is called write-back if the writes are done at a later time. Write-back is more efficient than write-through, but also a bit more prone to errors: if the machine crashes, or the power is cut at a bad moment, or the floppy is removed from the disk drive before the data in the cache waiting to be written gets written, the changes in the cache are usually lost. This might even mean that the filesystem (if there is one) is not in full working order, perhaps because the unwritten data held important changes to the bookkeeping information.

Because of this, you should never turn off the power without using a proper shutdown procedure (see chapter 6), or remove a floppy from the disk drive until it has been unmounted (if it was mounted) or after whatever program is using it has signaled that it is finished and the floppy drive light doesn’t shine anymore. The sync command flushes the buffer, i.e., forces all unwritten data to be written to disk, and can be used when one wants to be sure that everything is safely written. In traditional UNIX systems, there is a program called update running in the background which does a sync every 30 seconds, so it is usually not necessary to use sync. Linux has an additional daemon, bdflush, which does a more imperfect sync more frequently to avoid the sudden freeze due to heavy disk I/O that sync sometimes causes.

3Except a RAM disk, for obvious reasons.

Under Linux, bdflush is started by update. There is usually no reason to worry about it, but if bdflush happens to die for some reason, the kernel will warn about this, and you should start it by hand (/sbin/update).

The cache does not actually buffer files, but blocks, which are the smallest units of disk I/O (under Linux, they are usually 1 kB). This way, also directories, super blocks, other filesystem bookkeeping data, and non-filesystem disks are cached.

The effectiveness of a cache is primarily decided by its size. A small cache is next to useless: it will hold so little data that all cached data is flushed from the cache before it is reused. The critical size depends on how much data is read and written, and how often the same data is accessed. The only way to know is to experiment.

If the cache is of a fixed size, it is not very good to have it too big, either, because that might make the free memory too small and cause swapping (which is also slow). To make the most efficient use of real memory, Linux automatically uses all free RAM for buffer cache, but also automatically makes the cache smaller when programs need more memory.

Under Linux, you do not need to do anything to make use of the cache, it happens completely automatically. Except for following the proper procedures for shutdown and removing floppies, you do not need to worry about it.

Boots And Shutdowns

Start me up

Ah… you’ve got to… you’ve got to

Never, never never stop

Start it up

Ah… start it up, never, never, never You make a grown man cry, you make a grown man cry

(Rolling Stones)

This section explains what goes on when a Linux system is brought up and taken down, and how it should be done properly. If proper procedures are not followed, files might be corrupted or lost.

6.1 An overview of boots and shutdowns

The act of turning on a computer system and causing its operating system to be loaded^{1} is called booting. The name comes from an image of the computer pulling itself up from its bootstraps, but the act itself slightly more realistic.

During bootstrapping, the computer first loads a small piece of code called the bootstrap loader, which in turn loads and starts the operating system. The bootstrap loader is usually stored in a fixed location on a hard disk or a floppy. The reason for this two step process is that the operating system is big and complicated, but the first piece of code that the computer loads must be very small (a few hundred bytes), to avoid making the firmware unnecessarily complicated.

1On early computers, it wasn’t enough to merely turn on the computer, you had to manually load the operating system as well. These new-fangled thing-a-ma-jigs do it all by themselves.

Different computers do the bootstrapping differently. For PC’s, the computer (its BIOS) reads in the first sector (called the boot sector) of a floppy or hard disk. The bootstrap loader is contained within this sector. It loads the operating system from elsewhere on the disk (or from some other place).

After Linux has been loaded, it initializes the hardware and device drivers, and then runs init. init starts other processes to allow users to log in, and do things. The details of this part will be discussed below.

In order to shut down a Linux system, first all processes are told to terminate (this makes them close any files and do other necessary things to keep things tidy), then filesystems and swap areas are unmounted, and finally a message is printed to the console that the power can be turned off. If the proper procedure is not followed, terrible things can and will happen; most importantly, the filesystem buffer cache might not be flushed, which means that all data in it is lost and the filesystem on disk is inconsistent, and therefore possibly unusable.

6.2 The boot process in closer look

You can boot Linux either from a floppy or from the hard disk. The installation section in the Installation and Getting Started guide ([Wel]) tells you how to install Linux so you can boot it the way you want to.

When a PC is booted, the BIOS will do various tests to check that everything looks all right,^{2} and will then start the actual booting. It will choose a disk drive (typically the first floppy drive, if there is a floppy inserted, otherwise the first hard disk, if one is installed in the computer; the order might be configurable, however) and will then read its very first sector. This is called the boot sector; for a hard disk, it is also called the master boot record, since a hard disk can contain several partitions, each with their own boot sectors.

The boot sector contains a small program (small enough to fit into one sector) whose responsibility is to read the actual operating system from the disk and start it. When booting Linux from a floppy disk, the boot sector contains code that just reads the first few hundred blocks (depending on the actual kernel size, of course) to a predetermined place in memory. On a Linux boot floppy, there is no filesystem, the kernel is just stored in consecutive sectors, since this simplifies the boot process. It is possible, however, to boot from a floppy with a filesystem, by using LILO, the LInux LOader.

2This is called the power on self test, or POST for short.

When booting from the hard disk, the code in the master boot record will examine the partition table (also in the master boot record), identify the active partition (the partition that is marked to be bootable), read the boot sector from that partition, and then start the code in that boot sector. The code in the partition’s boot sector does what a floppy disk’s boot sector does: it will read in the kernel from the partition and start it. The details vary, however, since it is generally not useful to have a separate partition for just the kernel image, so the code in the partition’s boot sector can’t just read the disk in sequential order, it has to find the sectors wherever the filesystem has put them. There are several ways around this problem, but the most common way is to use LILO. (The details about how to do this are irrelevant for this discussion, however; see the LILO documentation for more information; it is most thorough.)

When booting with LILO, it will normally go right ahead and read in and boot the default kernel. It is also possible to configure LILO to be able to boot one of several kernels, or even other operating systems than Linux, and it is possible for the user to choose which kernel or operating system is to be booted at boot time. LILO can be configured so that if one holds down the alt , shift , or ctrl key at boot time (when LILO is loaded), LILO will ask what is to be booted and not boot the default right away. Alternatively, LILO can be configured so that it will always ask, with an optional timeout that will cause the default kernel to be booted.

With LILO, it is also possible to give a kernel command line argument, after the name of the kernel or operating system.

META: The are other boot loaders than LILO. Information about them will be added in some future version. loadlin.

Booting from floppy and from hard disk have both their advantages, but generally booting from the hard disk is nicer, since it avoids the hassle of playing around with floppies. It is also faster. However, it can be more troublesome to install the system to boot from the hard disk, so many people will first boot from floppy, then, when the system is otherwise installed and working well, will install LILO and start booting from the hard disk.

After the Linux kernel has been read into the memory, by whatever means, and is started for real, roughly the following things happen:

• The Linux kernel is installed compressed, so it will first uncompress itself. The

beginning of the kernel image contains a small program that does this.

• If you have a super-VGA card that Linux recognizes and that has some special text modes (such as 100 columns by 40 rows), Linux asks you which mode you want to use. During the kernel compilation, it is possible to preset a video mode, so that this is never asked. This can also be done with LILO or rdev.

• After this, the kernel checks what other hardware there is (hard disks, floppies, network adapters…), and configures some of its device drivers appropriately; while it does this, it outputs messages about its findings. For example, when I boot, I it looks like this:

     LILO boot:
     Loading linux.
     Console: colour EGA+ 80x25, 8 virtual consoles
     Serial driver version 3.94 with no serial options enabled
     tty00 at 0x03f8 (irq = 4) is a 16450
     tty01 at 0x02f8 (irq = 3) is a 16450
     lp_init: lp1 exists (0), using polling driver
     Memory: 7332k/8192k available (300k kernel code, 384k reserved, 176k data)
     Floppy drive(s): fd0 is 1.44M, fd1 is 1.2M
     Loopback device init
     Warning WD8013 board not found at i/o = 280.
     Math coprocessor using irq13 error reporting.
     Partition check:
       hda: hda1 hda2 hda3
     VFS: Mounted root (ext filesystem).
     Linux version 0.99.pl9-1 (root@haven) 05/01/93 14:12:20

The exact texts are different on different systems, depending on the hardware, the version of Linux being used, and how it has been configured.

• Then the kernel will try to mount the root filesystem. The place is configurable at compilation time, or any time with rdev or LILO. The filesystem type is detected automatically. If the mounting of the root filesystem fails, for example because you didn’t remember to include the corresponding filesystem driver in the kernel, the kernel panics and halts the system (there isn’t much it can do, anyway).

The root filesystem is usually mounted read-only (this can be set in the same way as the place). This makes it possible to check the filesystem while it is mounted; it is not a good idea to check a filesystem that is mounted read-write.

• After this, the kernel starts the program init (located in /sbin/init) in the background (this will always become process number 1). init does various startup chores. The exact things it does depends on how it is configured; see chapter 7 for more information (not yet written). It will at least start some essential background daemons.

• init then switches to multi-user mode, and starts a getty for virtual consoles and serial lines. getty is the program which lets people log in via virtual consoles and serial terminals. init may also start some other programs, depending on how it is configured.

• After this, the boot is complete, and the system is up and running normally.

6.3 More about shutdowns

It is important to follow the correct procedures when you shut down a Linux system. If you fail do so, your filesystems probably will become trashed and the files probably will become scrambled. This is because Linux has a disk cache that won’t write things to disk at once, but only at intervals. This greatly improves performance but also means that if you just turn off the power at a whim the cache may hold a lot of data and that what is on the disk may not be a fully working filesystem (because only some things have been written to the disk).

Another reason against just flipping the power switch is that in a multi-tasking system there can be lots of things going on in the background, and shutting the power can be quite disastrous. By using the proper shutdown sequence, you ensure that all background processes can save their data.

The command for properly shutting down a Linux system is shutdown. It is usually used in one of two ways.

If you are running a system where you are the only user, the usual way of using shutdown is to quit all running programs, log out on all virtual consoles, log in as root on one of them (or stay logged in as root if you already are, but you should change to the root directory, to avoid problems with unmounting), then give the command shutdown -h now (substitute now with a plus sign and a number in minutes if you want a delay, though you usually don’t on a single user system).

Alternatively, if your system has many users, use the command shutdown -h +time message, where time is the time in minutes until the system is halted, and message is a short explanation of why the system is shutting down.

     # shutdown -h +10 'We will install a new disk.  System should

     > be back on-line in three hours.'
     #

This will warn everybody that the system will shut down in ten minutes, and that they’d better get lost or lose data. The warning is printed to every terminal on which someone is logged in, including all xterms:

     Broadcast message from root (ttyp0) Wed Aug 2 01:03:25 1995...
     We will install a new disk.  System should
     be back on-line in three hours.
     The system is going DOWN for system halt in 10 minutes !!

The warning is automatically repeated a few times before the boot, with shorter and shorter intervals as the time runs out.

When the real shutting down starts after any delays, all filesystems (except the root one) are unmounted, user processes (if anybody is still logged in) are killed, daemons are shut down, all filesystem are unmounted, and generally everything settles down. When that is done, init prints out a message that you can power down the machine. Then, and only then, should you move your fingers towards the power switch.

Sometimes, although rarely on any good system, it is impossible to shut down properly. For instance, if the kernel panics and crashes and burns and generally misbehaves, it might be completely impossible to give any new commands, hence shutting down properly is somewhat difficult, and just about everything you can do is hope that nothing has been too severely damaged and turn off the power. If the troubles are a bit less severe (say, somebody hit your keyboard with an axe), and the kernel and the update program still run normally, it is probably a good idea to wait a couple of minutes to give update a chance to flush the buffer cache, and only cut the power after that.

Some people like to shut down using the command sync^{3} three times, waiting for the disk I/O to stop, then turn off the power. If there are no running programs, this is about equivalent to using shutdown. However, it does not unmount any filesystems and this can lead to problems with the ext2fs “clean filesystem” flag. The triple-sync method is not recommended.

(In case you’re wondering: the reason for three syncs is that in the early days of UNIX, when the commands were typed separately, that usually gave sufficient time for most disk I/O to be finished.)

3sync flushes the buffer cache.

6.4 Rebooting

Rebooting means booting the system again. This can be accomplished by first shutting it down completely, turning power off, and then turning it back on. A simpler way is to ask shutdown to reboot the system, instead of merely halting it. This is accomplished by using the -r option to shutdown, for example, by giving the command

shutdown -r now.

Most Linux systems run shutdown -r now when ctrl-alt-del is pressed on the keyboard. This reboots the system. The action on ctrl-alt-del is configurable, however, and it might be better to allow for some delay before the reboot on a multiuser machine. Systems that are physically accessible to anyone might even be configured to do nothing when ctrl-alt-del is pressed.

6.5 Single user mode

The shutdown command can also be used to bring the system down to single user mode, in which no one can log in, but root can use the console. This is useful for system administration tasks that can’t be done while the system is running normally. Single user mode is discussed more thoroughly in chapter ??.

6.6 Emergency boot floppies

It is not always possible to boot a computer from the hard disk. For example, if you make a mistake in configuring LILO, you might make your system unbootable. For these situations, you need an alternative way of booting that will always work (as long as the hardware works). For typical PC’s, this means booting from the floppy drive.

Most Linux distributions allow one to create an emergency boot floppy during installation. It is a good idea to do this. However, some such boot disks contain only the kernel, and assume you will be using the programs on the distribution’s installation disks to fix whatever problem you have. Sometimes those programs aren’t enough; for example, you might have to restore some files from backups made with software not on the installation disks.

Thus, it might be necessary to create a custom root floppy as well. The Bootdisk HOWTO by Graham Chapman ([Cha]) contains instructions for doing this. You must, of course, remember to keep your emergency boot and root floppies up to date.

You can’t use the floppy drive you use to mount the root floppy for anything else. This can be inconvenient if you only have one floppy drive. However, if you have enough memory, you can configure your boot floppy to load the root disk to a ramdisk (the boot floppy’s kernel needs to be specially configured for this). Once the root floppy has been loaded into the ramdisk, the floppy drive is free to mount other disks.

init

Uuno on numero yksi

This chapter describes the init process, which is the first user level process started by the kernel. init has many important duties, such as starting getty (so that users can log in), implementing run levels, and taking care of orphaned processes. This chapter explains how init is configured and how you can make use of the different run levels.

7.1 init comes first

init is one of those programs that are absolutely essential to the operation of a Linux system, but that you still can mostly ignore. A good Linux distribution will come with a configuration for init that will work for most systems, and on these systems there is nothing you need to do about init. Usually, you only need to worry about init if you hook up serial terminals, dial-in (not dial-out) modems, or if you want to change the default run level.

When the kernel has started itself (has been loaded into memory, has started running, and has initialized all device drivers and data structures and such), it finishes its own part of the boot process by starting a user level program, init. Thus, init is always the first process (its process number is always 1).

The kernel looks for init in a few locations that have been historically used for it, but the proper location for it (on a Linux system) is /sbin/init. If the kernel can’t find init, it tries to run /bin/sh, and if that also fails, the startup of the system fails.

When init starts, it finishes the boot process by doing a number of administrative tasks, such as checking filesystems, cleaning up /tmp, starting various services, and starting a getty for each terminal and virtual console where users should be able to log in (see chapter 8).

After the system is properly up, init restarts getty for each terminal after a user has logged out (so that the next user can log in). init also adopts orphan processes: when a process starts a child process and dies before its child, the child immediately becomes a child of init. This is important for various technical reasons, but it is good to know it, since it makes it easier to understand process lists and process tree graphs.^{1}

There are a few variants of init available. Most Linux distributions use sysvinit (written by Miquel van Smoorenburg), which is based on the System V init design. The BSD versions of Unix have a different init. The primary difference is run levels: System V has them, BSD does not (at least traditionally). This difference is not essential. We’ll look at sysvinit only.

7.2 Configuring init to start getty: the /etc/inittab file

When it starts up, init reads the /etc/inittab configuration file. While the system is running, it will re-read it, if sent the HUP signal;^{2} this feature makes it unnecessary to boot the system to make changes to the init configuration take effect.

The /etc/inittab file is a bit complicated. We’ll start with the simple case of configuring getty lines. Lines in /etc/inittab consist of four colon-delimited fields:

id:runlevels:action:process

The fields are described below. In addition, /etc/inittab can contain empty lines, and lines that begin with a number sign (‘#’); these are both ignored.

id This identifies the line in the file. For getty lines, it specifies the terminal it runs on (the characters after /dev/tty in the device file name). For other lines, it doesn’t matter (except for length restrictions), but it should be unique. runlevels The run levels the line should be considered for. The run levels are

1init itself is not allowed to die. You can’t kill init even with SIGKILL.

2Using the command kill -HUP 1 as root, for example

given as single digits, without delimiters. (Run levels are described in the next section.)

action What action should be taken by the line, e.g., respawn to run the command in the next field again, when it exits, or once to run it just once.

process The command to run.

To start a getty on the first virtual terminal (/dev/tty1), in all the normal multiuser run levels (2–5), one would write the following line:

     1:2345:respawn:/sbin/getty 9600 tty1

The first field says that this is the line for /dev/tty1. The second field says that it applies to run levels 2, 3, 4, and 5. The third field means that the command should be run again, after it exits (so that one can log in, log out, and then log in again). The last field is the command that runs getty on the first virtual terminal.^{3}

If you wanted to add terminals or dial-in modem lines to a system, you’d add more lines to /etc/inittab, one for each terminal or dial-in line. For more details, see the manual pages init(8), inittab(5), and getty(8).

If a command fails when it starts, and init is configured to restart it, it will use a lot of system resources: init starts it, it fails, init starts it, it fails, init starts it, it fails, and so on, ad infinitum. To prevent this, init will keep track of how often it restarts a command, and if the frequency grows to high, it will delay for five minutes before restarting again.

7.3 Run levels

A run level is a state of init and the whole system that defines what system services are operating. Run levels are identified by numbers, see table 7.1. There is no consensus of how to use the user defined run levels (2 through 5). Some system administrators use run levels to define which subsystems are working, e.g., whether X is running, whether the network is operational, and so on. Others have all subsystems always running or start and stop them individually, without changing run levels, since run levels are too coarse for controlling their systems. You need to decide for yourself, but it might be easiest to follow the way your Linux distribution does things.

3Different versions of getty are run differently. Consult your manual page—and make sure it is the correct manual page.

Table 7.1: Run level numbers

0  Halt the system.
1  Single-user mode (for special administration).
2–5 Normal operation (user defined).
6  Reboot.

Run levels are configured in /etc/inittab by lines like the following:

     l2:2:wait:/etc/init.d/rc 2

The first field is an arbitrary label, the second one means that this applies for run level 2. The third field means that init should run the command in the fourth field once, when the run level is entered, and that init should wait for it to complete. The /etc/init.d/rc command runs whatever commands are necessary to start and stop services to enter run level 2.

The command in the fourth field does all the hard work of setting up a run level. It starts services that aren’t already running, and stops services that shouldn’t be running in the new run level any more. Exactly what the command is, and how run levels are configured, depends on the Linux distribution.

When init starts, it looks for a line in /etc/inittab that specifies the default run level:

     id:2:initdefault:

You can ask init to go to a non-default run level at startup by giving the kernel a command line argument of single or emergency.^{4} This allows you to choose the single user mode (run level 1), which is described in section 7.5.

While the system is running, the telinit command can change the run level. When the run level is changed, init runs the relevant command from /etc/inittab.

7.4 Special configuration in /etc/inittab

The /etc/inittab has some special features that allow init to react to special circumstances. These special features are marked by special keywords in the third field. Some examples:

4Kernel command line arguments can be given via LILO, for example. See section 7.5.

powerwait Allows init to shut the system down, when the power fails. This assumes the use of a UPS, and software that watches the UPS and informs init that the power is off.

ctrlaltdel Allows init to reboot the system, when the user presses control-alt-del on the console keyboard. Note that the system administrator can configure the reaction to C-A-D to be something else instead, e.g., to be ignored, if the system is in a public location.^{5}

sysinit Command to be run when the system is booted. This command usually cleans up /tmp, for example.

The list above is not exhaustive. See your inittab(5) manual page for all possibilities, and for details on how to use the above ones.

7.5 Booting in single user mode

An important run level is single user mode (run level 1), in which only the system administrator is using the machine and as few system services—including logins—as possible are running. Single user mode is necessary for a few administrative tasks,^{6} such as running fsck on a /usr partition—this requires that the partition be unmounted, and that can’t happen, unless just about all system services are killed.

A running system can be taken to single user mode by using telinit to request run level 1. At bootup, it can be entered by giving the word single or emergency on the kernel command line: the kernel gives the command line to init as well, and init understands from that word that it shouldn’t use the default run level. (The kernel command line is entered in a way that depends on how you boot the system.)

Booting into single user mode is sometimes necessary so that one can run fsck by hand, before anything mounts or otherwise touches a broken /usr partition (any activity on a broken filesystem is likely to break it more, so fsck should be run as soon as possible).

The bootup scripts init runs will automatically enter single user mode, if the automatic fsck at bootup fails. This is an attempt to prevent the system from using a filesystem that is so broken that fsck can’t fix it automatically. Such breakage is relatively rare, and usually involves a broken hard disk or an experimental kernel release, but it’s good to be prepared.

5Or to start nethack.

6It probably shouldn’t be used for playing nethack.

As a security measure, a properly configured system will ask for the root password before starting the shell in single user mode. Otherwise, it would be simple to just enter a suitable line to LILO to get in as root. (This will break if /etc/passwd has been broken by filesystem problems, of course, and in that case you’d better have a boot floppy handy.)

Logging In And Out

This chapter needs a quote. Suggestions, anyone?

This section describes what happens when a user logs in or out. The various interactions of background processes, log files, configuration files, and so on are described in some detail.

8.1 Logins via terminals

Figure 8.1 shows how logins happen via terminals. First, init makes sure there is a getty program for the terminal connection (or console). getty listens at the terminal and waits for the user to notify that he is ready to login in (this usually means that the user must type something). When it notices a user, getty outputs a welcome message (stored in /etc/issue), and prompts for the username, and finally runs the login program. login gets the username as a parameter, and prompts the user for the password. If these match, login starts the shell configured for the user; else it just exits and terminates the process (perhaps after giving the user another chance at entering the username and password). init notices that the process terminated, and starts a new getty for the terminal.

Note that the only new process is the one created by init (using the fork system call); getty and login only replace the program running in the process (using the exec system call).

A separate program, for noticing the user, is needed for serial lines, since it can be (and traditionally was) complicated to notice when a terminal becomes active. getty also adapts to the speed and other settings of the connection, which is important especially for dial-in connections, where these parameters may change from call to call.

There are several versions of getty and init in use, all with their good and bad points. It is a good idea to learn about the versions on your system, and also about the other versions (you could use the Linux Software Map to search them). If you don’t have dial-in’s, you probably don’t have to worry about getty, but init is still important.

8.2 Logins via the network

Two computers in the same network are usually linked via a single physical cable. When they communicate over the network, the programs in each computer that take part in the communication are linked via a virtual connection, a sort of imaginary cable. As far as the programs at either end of the virtual connection are concerned, they have a monopoly on their own cable. However, since the cable is not real, only imaginary, the operating systems of both computers can have several virtual connections share the same physical cable. This way, using just a single cable, several programs can communicate without having to know of or care about the other communications. It is even possible to have several computers use the same cable; the virtual connections exist between two computers, and the other computers ignore those connections that they don’t take part in.

That’s a complicated and over-abstracted description of the reality. It might, however, be good enough to understand the important reason why network logins are somewhat different from normal logins. The virtual connections are established when there are two programs on different computers that wish to communicate. Since it is in principle possible to login from any computer in a network to any other computer, there is a huge number of potential virtual communications. Because of this, it is not practical to start a getty for each potential login.

There is a single process inetd (corresponding to getty) that handles all network logins. When it notices an incoming network login (i.e., it notices that it gets a new virtual connection to some other computer), it starts a new process to handle that single login. The original process remains and continues to listen for new logins.

To make things a bit more complicated, there is more than one communication protocol for network logins. The two most important ones are telnet and rlogin. In addition to logins, there are many other virtual connections that may be made (for FTP, Gopher, HTTP, and other network services). It would be ineffective to have a separate process listening for a particular type of connection, so instead there is only one listener that can recognize the type of the connection and can start the correct type of program to provide the service. This single listener is called inetd; see the “Linux Network Administrators’ Guide” for more information.

8.3 What login does

The login program takes care of authenticating the user (making sure that the username and password match), and of setting up an initial environment for the user by setting permissions for the serial line and starting the shell.

Part of the initial setup is outputting the contents of the file /etc/motd (short for message of the day) and checking for electronic mail. These can be disabled by creating a file called .hushlogin in the user’s home directory.

If the file /etc/nologin exists, logins are disabled. That file is typically created by shutdown and relatives. login checks for this file, and will refuse to accept a login if it exists. If it does exist, login outputs its contents to the terminal before it quits.

login logs all failed login attempts in a system log file (via syslog). It also logs all logins by root. Both of these can be useful when tracking down intruders.

Currently logged in people are listed in /var/run/utmp. This file is valid only until the system is next rebooted or shut down; it is cleared when the system is booted. It lists each user and the terminal (or network connection) he is using, along with some other useful information. The who, w, and other similar commands look in utmp to see who are logged in.

All successful logins are recorded into /var/log/wtmp. This file will grow without limit, so it must be cleaned regularly, for example by having a weekly cron job to clear it.^{1} The last command browses wtmp.

Both utmp and wtmp are in a binary format (see the utmp manual page); it is unfortunately not convenient to examine them without special programs.

8.4 X and xdm

META: X implements logins via xdm; also: xterm -ls

1Good Linux distributions do this out of the box.

8.5 Access control

The user database is traditionally contained in the /etc/passwd file. Some systems use shadow passwords, and have moved the passwords to /etc/shadow. Sites with many computers that share the accounts use NIS or some other method to store the user database; they might also automatically copy the database from one central location to all other computers.

The user database contains not only the passwords, but also some additional information about the users, such as their real names, home directories, and login shells. This other information needs to be public, so that anyone can read it. Therefore the password is stored encrypted. This does have the drawback that anyone with access to the encrypted password can use various cryptographical methods to guess it, without trying to actually log into the computer. Shadow passwords try to avoid this by moving the password into another file, which only root can read (the password is still stored encrypted). However, installing shadow passwords later onto a system that did not support them can be difficult.

With or without passwords, it is important to make sure that all passwords in a system are good, i.e., not easily guessable. The crack program can be used to crack passwords; any password it can find is by definition not a good one. While crack can be run by intruders, it can also be run by the system adminstrator to avoid bad passwords. Good passwords can also be enforced by the passwd program; this is in fact more effective in CPU cycles, since cracking passwords requires quite a lot of computation.

The user group database is kept in /etc/group; for systems with shadow passwords, there can be a /etc/shadow.group.

root usually can’t login via most terminals or the network, only via terminals listed in the /etc/securetty file. This makes it necessary to get physical access to one of these terminals. It is, however, possible to log in via any terminal as any other user, and use the su command to become root.

8.6 Shell startup

When an interactive login shell starts, it automatically executes one or more predefined files. Different shells execute different files; see the documentation of each shell for further information.

Most shells first run some global file, for example, the Bourne shell (/bin/sh) and its derivatives execute /etc/profile; in addition, they execute .profile in the user’s home directory. /etc/profile allows the system administrator to have set up a common user environment, especially by setting the PATH to include local command directories in addition to the normal ones. On the other hand, .profile allows the user to customize the environment to his own tastes by overriding, if necessary, the default environment.

8.6. Shell startup 83 Figure 8.1: Logins via terminals: the interaction of init, getty, login, and the shell.

Managing user accounts

The similarities of sysadmins and drug dealers: both measure stuff in K’s, and both have users.

(Old, tired computer joke.)

This chapter explains how to create new user accounts, how to modify the properties of those accounts, and how to remove the accounts. Different Linux systems have different tools for doing this.

9.1 What’s an account?

When a computer is used by many people it is usually necessary to differentiate between the users, for example, so that their private files can be kept private. This is important even if the computer can only be used by a single person at a time, as with most microcomputers.^{1} Thus, each user is given a unique username, and that name is used to log in.

There’s more to a user than just a name, however. An account is all the files, resources, and information belonging to one user. The term hints at banks, and in a commercial system each account usually has some money attached to it, and that money vanishes at different speeds depending on how much the user stresses the system. For example, disk space might have a price per megabyte and day, and processing time might have a price per second.

1It might be quite embarrassing if my sister could read my love letters.

9.2 Creating a user

The Linux kernel itself treats users are mere numbers. Each user is identified by a unique integer, the user id or uid, because numbers are faster and easier for a computer to process than textual names. A separate database outside the kernel assigns a textual name, the username, to each user id. The database contains additional information as well.

To create a user, you need to add information about the user to the user database, and create a home directory for him. It may also be necessary to educate the user, and set up a suitable initial environment for him.

Most Linux distributions come with a program for creating accounts. There are several such programs available. Two command line alternatives are adduser and useradd; there may be a GUI tool as well. Whatever the program, the result is that there is little if any manual work to be done. Even if the details are many and intricate, these programs make everything seem trivial. However, section 9.2.4 describes how to do it by hand.

9.2.1 /etc/passwd and other informative files

The basic user database in a Unix system is the text file, /etc/passwd (called the password file), which lists all valid usernames and their associated information. The file has one line per username, and is divided into seven colon-delimited fields:

1. Username.

2. Password, in an encrypted form.

3. Numeric user id.

4. Numeric group id.

5. Full name or other description of account.

6. Home directory.

7. Login shell (program to run at login).

The format is explained in more detail in passwd(5).

Any user on the system may read the password file, so that they can, for example, learn the name of another user. This means that the password (the second field) is also available to everyone. The password file encrypts the password, so in theory there is no problem. However, the encryption is breakable, especially if the password is weak (e.g., it is short or it can be found in a dictionary). Therefore it is not a good idea to have the password in the password file.

Many Linux systems have shadow passwords. This is an alternative way of storing the password: the encrypted password is stored in a separate file, /etc/shadow, which only root can read. The /etc/passwd file only contains a special marker in the second field. Any program that needs to verify a user is setuid, and can therefore access the shadow password file. Normal programs, which only use the other fields in the password file, can’t get at the password.^{2}

9.2.2 Picking numeric user and group ids

On most systems it doesn’t matter what the numeric user and group ids are, but if you use the Network filesystem (NFS), you need to have the same uid and gid on all systems. This is because NFS also identifies users with the numeric uids. If you aren’t using NFS, you can let your account creation tool pick them automatically.

If you are using NFS, you’ll have to be invent a mechanism for synchronizing account information. One alternative is to the NIS system (see [Kir]).

META: this is wrong place? However, you should try to avoid re-using numeric uid’s (and textual usernames), because the new owner of the uid (or username) may get access to the old owner’s files (or mail, or whatever).

9.2.3 Initial environment: /etc/skel

When the home directory for a new user is created, it is initialized with files from the /etc/skel directory. The system administrator can create files in /etc/skel that will provide a nice default environment for users. For example, he might create a /etc/skel/.profile that sets the EDITOR environment variable to some editor that is friendly towards new users.

However, it is usually best to try to keep /etc/skel as small as possible, since it will be next to impossible to update existing users’ files. For example, if the name of the friendly editor changes, all existing users would have to edit their .profile. The system administrator could try to do it automatically, with a script, but that is almost certain going to break someone’s file.

2Yes, this means that the password file has all the information about a user except his password. The wonder of development.

Whenever possible, it is better to put global configuration into global files, such as /etc/profile. This way it is possible to update it without breaking users’ own setups.

9.2.4 Creating a user by hand

To create a new account manually, follow these steps:

1. Edit /etc/passwd with vipw(8) and add a new line for the new account. Be careful with the syntax. Do not edit directly with an editor! vipw locks the file, so that other commands won’t try to update it at the same time. You should make the password field be ‘*’, so that it is impossible to log in.

2. Similarly, edit /etc/group with vigr, if you need to create a new group as well.

3. Create the home directory of the user with mkdir.

4. Copy the files from /etc/skel to the new home directory.

5. Fix ownerships and permissions with chown and chmod. The -R option is most useful. The correct permissions vary a little from one site to another, but usually the following commands do the right thing:

          cd /home/newusername
          chown -R username.group .
          chmod -R go=u,go-w .
          chmod go= .

6. Set the password with passwd(1).

After you set the password in the last step, the account will work. You shouldn’t set it until everything else has been done, otherwise the user may inadvertently log in while you’re still copying the files.

It is sometimes necessary to create dummy accounts^{3} that are not used by people. For example, to set up an anonymous FTP server (so that anyone can download files from it, without having to get an account first), you need to create an account called ftp. In such cases, it is usually not necessary to set the password (last step above). Indeed, it is better not to, so that no-one can use the account, unless they first become root, since root can become any user.

3Surreal users?

9.3 Changing user properties

There are a few commands for changing various properties of an account (i.e., the relevant field in /etc/passwd):

chfn Change the full name field.

chsh Change the login shell.

passwd Change the password.

The super-user may use these commands to change the properties of any account. Normal users can only change the properties of their own account. It may sometimes be necessary to disable these commands (with chmod) for normal users, for example in an environment with many novice users.

Other tasks need to be done by hand. For example, to change the username, you need to edit /etc/passwd directly (with vipw, remember). Likewise, to add or remove the user to more groups, you need to edit /etc/group (with vigr). Such tasks tend to be rare, however, and should be done with caution: for example, if you change the username, e-mail will no longer reach the user, unless you also create a mail alias.^{4}

9.4 Removing a user

To remove a user, you first remove all his files, mailboxes, mail aliases, print jobs, cron and at jobs, and all other references to the user. Then you remove the relevant lines from /etc/passwd and /etc/group (remember to remove the username from all groups it’s been added to). It may be a good idea to first disable the account (see below), before you start removing stuff, to prevent the user from using the account while it is being removed.

Remember that users may have files outside their home directory. The find command can find them:

     find / -user username

However, note that the above command will take a long time, if you have large disks. If you mount network disks (see section 2.3.8), you need to be careful so that you won’t trash the network or the server.

4The user’s name might change due to marriage, for example, and he might want to have his username reflect his new name.

Some Linux distributions come with special commands to do this; look for deluser or userdel. However, it is easy to do it by hand as well, and the commands might not do everything.

9.5 Disabling a user temporarily

It is sometimes necessary to temporarily disable an account, without removing it. For example, the user might not have paid his fees, or the system administrator may suspect that a cracker has got the password of that account.

The best way to disable an account is to change its shell into a special program that just prints a message. This way, whoever tries to log into the account, will fail, and will know why. The message can tell the user to contact the system administrator so that any problems may be dealt with.

It would also be possible to change the username or password to something else, but then the user won’t know what is going on. Confused users mean more work.^{5}

A simple way to create the special programs is to write ‘tail scripts’:

    #!/usr/bin/tail +2
    This account has been closed due to a security breach.
    Please call 555-1234 and wait for the men in black to arrive.

The first two characters (‘#!’) tell the kernel that the rest of the line is a command that needs to be run to interpret this file. The tail command in this case outputs everything except the first line to the standard output.

If billg is suspected of a security breach, the system administrator would do something like this:

     # chsh -s /usr/local/lib/no-login/security billg
     # su - tester
     This account has been closed due to a security breach.
     Please call 555-1234 and wait for the men in black to arrive.
     #

The purpose of the su is to test that the change worked, of course.

Tail scripts should be kept in a separate directory, so that their names don’t interfere with normal user commands.

5But they can be so fun, if you’re a BOFH.

Backups

Hardware is indeterministically reliable. Software is deterministically unreliable. People are indeterministically unreliable.

Nature is deterministically reliable.

This chapter explains about why, how, and when to make backups, and how to restore things from backups.

10.1 On the importance of being backed up

Your data is valuable. It will cost you time and effort re-create it, and that costs money or at least personal grief and tears; sometimes it can’t even be re-created, e.g., if it is the results of some experiments. Since it is an investment, you should protect it and take steps to avoid losing it.

There are basically four reasons why you might lose data: hardware failures, software bugs, human action, or natural disasters.^{1} Although modern hardware tends to be quite reliable, it can still break seemingly spontaneously. The most critical piece of hardware for storing data is the hard disk, which relies on tiny magnetic fields remaining intact in a world filled with electromagnetic noise. Modern software doesn’t even tend to be reliable; a rock solid program is an exception, not a rule. Humans are quite unreliable, they will either make a mistake, or they will be malicious and destroy data on purpose. Nature might not be evil, but it can wreak havoc even when being good. All in all, it is a small miracle that anything works at all.

1The fifth reason is “something else”.

Backups are a way to protect the investment in data. By having several copies of the data, it does not matter as much if one is destroyed (the cost is only that of the restoration of the lost data from the backup).

It is important to do backups properly. Like everything else that is related to the physical world, backups will fail sooner or later. Part of doing backups well is to make sure they work; you don’t want to notice that your backups didn’t work.^{2} Adding insult to injury, you might have a bad crash just as you’re making the backup; if you have only one backup medium, it might destroyed as well, leaving you with the smoking ashes of hard work.^{3} Or you might notice, when trying to restore, that you forgot to back up something important, like the user database on a 15 000 user site. Best of all, all your backups might be working perfectly, but the last known tape drive reading the kind of tapes you used was the one that now has a bucketful of water in it.

When it comes to backups, paranoia is in the job description.

10.2 Selecting the backup medium

The most important decision regarding backups is the choice of backup medium. You need to consider cost, reliability, speed, availability, and usability.

Cost is important, since you should preferably have several times more backup storage than what you need for the data. A cheap medium is usually a must.

Reliability is extremely important, since a broken backup can make a grown man cry. A backup medium must be able to hold data without corruption for years. The way you use the medium affects it reliability as a backup medium. A hard disk is typically very reliable, but as a backup medium it is not very reliable, if it is in the same computer as the disk you are backing up.

Speed is usually not very important, if backups can be done without interaction. It doesn’t matter if a backup takes two hours, as long as it needs no supervision. On the other hand, if the backup can’t be done when the computer would otherwise be idle, then speed is an issue.

Availability is obviously necessary, since you can’t use a backup medium if it doesn’t exist. Less obvious is the need for the medium to be available even in the future, and on computers other than your own. Otherwise you may not be able to restore your backups after a disaster.

2Don’t laugh. This has happened to several people.

3Been there, done that…

Usability is a large factor in how often backups are made. The easier it is to make backups, the better. A backup medium mustn’t be hard or boring to use.

The typical alternatives are floppies and tapes. Floppies are very cheap, fairly reliable, not very fast, very available, but not very usable for large amounts of data. Tapes are cheap to somewhat expensive, fairly reliable, fairly fast, quite available, and—depending on the size of the tape—quite comfortable.

There are other alternatives. They are usually not very good on availability, but if that is not a problem, they can be better in other ways. For example, magneto-optical disks can have good sides of both floppies (they’re random access, making restoration of a single file quick) and tapes (contain a lot of data).

10.3 Selecting the backup tool

There are many tools that can be used to make backups. The traditional UNIX tools used for backups are tar, cpio, and dump. In addition, there are large number of third party packages (both freeware and commercial) that can be used. The choice of backup medium can affect the choice of tool.

tar and cpio are similar, and mostly equivalent from a backup point of view. Both are capable of storing files on tapes, and retrieving files from them. Both are capable of using almost any media, since the kernel device drivers take care of the low level device handling and the devices all tend to look alike to user level programs. Some UNIX versions of tar and cpio may have problems with unusual files (symbolic links, device files, files with very long pathnames, and so on), but the Linux versions should handle all files correctly.

dump is different in that it reads the filesystem directly and not via the filesystem. It is also written specifically for backups; tar and cpio are really for archiving files, although they work for backups as well.

Reading the filesystem directly has some advantages. It makes it possible to back files up without affecting their time stamps; for tar and cpio, you would have to mount the filesystem read-only first. Directly reading the filesystem is also more effective, if everything needs to be backed up, since it can be done with much less disk head movement. The major disadvantage is that it makes the backup program specific to one filesystem type; the Linux dump program understands the ext2 filesystem only.

dump also directly supports backup levels (which we’ll be discussing below); with tar and cpio this has to be implemented with other tools.

A comparison of the third party backup tools is beyond the scope of this book. The Linux Software Map lists many of the freeware ones.

10.4 Simple backups

A simple backup scheme is to back up everything once, then back up everything that has been modified since the previous backup. The first backup is called a full backup, the subsequent ones are incremental backups. A full backup is often more laborius than incremental ones, since there is more data to write to the tape and a full backup might not fit onto one tape (or floppy). Restoring from incremental backups can be many times more work than from a full one. Restoration can be optimized so that you always back up everything since the previous full backup; this way, backups are a bit more work, but there should never be a need to restore more than a full backup and an incremental backup.

If you want to make backups every day and have six tapes, you could use tape 1 for the first full backup (say, on a Friday), and tapes 2 to 5 for the incremental backups (Monday through Thursday). Then you make a new full backup on tape 6 (second Friday), and start doing incremental ones with tapes 2–5 again. You don’t want to overwrite tape 1 with until you’ve got a new full backup, lest something happens while you’re making the full backup. After you’ve made a full backup to tape 6, you want to keep tape 1 somewhere else, so that when your other backup tapes are destroyed in the fire, you still have at least something left. When you need to make the next full backup, you fetch tape 1 and leave tape 6 in its place.

If you have more than six tapes, you can use the extra ones for full backups. Each time you make a full backup, you use the oldest tape. This way you can have full backups from several previous weeks, which is good if you want to find an old, now deleted file, or an old version of a file.

10.4.1 Making backups with tar

A full backup can easily be made with tar:

     # tar –create –file /dev/ftape /usr/src
     tar:  Removing leading / from absolute path names in the archive

     #

The example above uses the GNU version of tar and its long option names. The traditional version of tar only understands single character options. The GNU version can also handle backups that don’t fit on one tape or floppy, and also very long paths; not all traditional versions can do these things. (Linux only uses GNU tar.)

If your backup doesn’t fit on one tape, you need to use the –multi-volume (-M) option:

     # tar -cMf /dev/fd0H1440 /usr/src
     tar:  Removing leading / from absolute path names in the archive
     Prepare volume #2 for /dev/fd0H1440 and hit return:
     #

Note that you should format the floppies before you begin the backup, or else use another window or virtual terminal and do it when tar asks for a new floppy.

After you’ve made a backup, you should check that it is OK, using the –compare (-d) option:

     # tar –compare –verbose -f /dev/ftape
     usr/src/
     usr/src/linux
     usr/src/linux-1.2.10-includes/
     ....
     #

Failing to check a backup means that you will not notice that your backups aren’t working until after you’ve lost the original data.

An incremental backup can be done with tar using the –newer (-N) option:

     # tar –create –newer '8 Sep 1995' –file /dev/ftape /usr/src –verbose
     tar:  Removing leading / from absolute path names in the archive
     usr/src/
     usr/src/linux-1.2.10-includes/
     usr/src/linux-1.2.10-includes/include/
     usr/src/linux-1.2.10-includes/include/linux/
     usr/src/linux-1.2.10-includes/include/linux/modules/
     usr/src/linux-1.2.10-includes/include/asm-generic/

10.4. Simple backups 95

usr/src/linux-1.2.10-includes/include/asm-i386/
usr/src/linux-1.2.10-includes/include/asm-mips/
usr/src/linux-1.2.10-includes/include/asm-alpha/
usr/src/linux-1.2.10-includes/include/asm-m68k/
usr/src/linux-1.2.10-includes/include/asm-sparc/
usr/src/patch-1.2.11.gz
#

Unfortunately, tar can’t notice when a file’s inode information has changed, for example, that it’s permission bits have been changed, or when its name has been changed. This can be worked around using find and comparing current filesystem state with lists of files that have been previously backed up. Scripts and programs for doing this can be found on Linux ftp sites.

10.4.2 Restoring files with tar

The –extract (-x) option for tar extracts files:

# tar –extract –same-permissions –verbose –file /dev/fd0H1440
usr/src/
usr/src/linux
usr/src/linux-1.2.10-includes/
usr/src/linux-1.2.10-includes/include/
usr/src/linux-1.2.10-includes/include/linux/
usr/src/linux-1.2.10-includes/include/linux/hdreg.h
usr/src/linux-1.2.10-includes/include/linux/kernel.h
...
#

You also extract only specific files or directories (which includes all their files and subdirectories) by naming on the command line:

# tar xpvf /dev/fd0H1440 usr/src/linux-1.2.10-includes/include/linux/hdreg.h
usr/src/linux-1.2.10-includes/include/linux/hdreg.h
#

Use the –list (-t) option, if you just want to see what files are on a backup volume:

# tar –list –file /dev/fd0H1440
usr/src/

     usr/src/linux
     usr/src/linux-1.2.10-includes/
     usr/src/linux-1.2.10-includes/include/
     usr/src/linux-1.2.10-includes/include/linux/
     usr/src/linux-1.2.10-includes/include/linux/hdreg.h
     usr/src/linux-1.2.10-includes/include/linux/kernel.h
     ...
     #

Note that tar always reads the backup volume sequentially, so for large volumes it is rather slow. It is not possible, however, to use random access database techniques when using a tape drive or some other sequential medium.

tar doesn’t handle deleted files properly. If you need to restore a filesystem from a full and an incremental backup, and you have deleted a file between the two backups, it will exist again after you have done the restore. This can be a big problem, if the file has sensitive data that should no longer be available.

10.5 Multilevel backups

The simple backup method outlined in the previous section is often quite adequate for personal use or small sites. For more heavy duty use, multilevel backups are more appropriate.

The simple method has two backup levels: full and incremental backups. This can be generalized to any number of levels. A full backup would be level 0, and the different levels of incremental backups levels 1, 2, 3, … At each incremental backup level you back up everything that has changed since the previous backup at the same or a previous level.

The purpose for doing this is that it allows a longer backup history cheaply. In the example in the previous section, the backup history went back to the previous full backup. This could be extended by having more tapes, but only a week per new tape, which might be too expensive. A longer backup history is useful, since deleted or corrupted files are often not noticed for a long time. Even a version of a file that is not very up to date is better than no file at all.

With multiple levels the backup history can be extended more cheaply. For example, if we buy ten tapes, we could use tapes 1 and 2 for monthly backups (first Friday each month), tapes 3 to 6 for weekly backups (other Fridays; note that there

Figure 10.1: A sample multilevel backup schedule.

can be five Fridays in one month, so we need four more tapes), and tapes 7 to 10 for daily backups (Monday to Thursday). With only four more tapes, we’ve been able to extend the backup history from two weeks (after all daily tapes have been used) to two months. It is true that we can’t restore every version of each file during those two months, but what we can restore is often good enough.

Figure 10.1 shows which backup level is used each day, and which backups can be restored from at the end of the month.

Backup levels can also be used to keep filesystem restoration time to a minimum. If you have many incremental backups with monotonously growing level numbers, you need to restore all of them if you need to rebuild the whole filesystem. Instead you can use level numbers that aren’t monotonous, and keep down the number of backups to restore.

To minimize the number of tapes needed to restore, you could use a smaller level for each incremental tape. However, then the time to make the backups increases (each backup copies everything since the previous full backup). A better scheme is suggested by the dump manual page and described by the table 10.2. Use the following succession of backup levels: 3, 2, 5, 4, 7, 6, 9, 8, 9… This keeps both the backup and restore times low. The most you have to backup is two day’s worth of work. The number of tapes for a restore depends on how long you keep between full backups, but it is less than in the simple schemes.

A fancy scheme can reduce the amount of labor needed, but it does mean there are more things to keep track of. You must decide if it is worth it.

dump has built-in support for backup levels. For tar and cpio it must be implemented with shell scripts.

Figure 10.2: Efficient backup scheme using many backup levels

Tape Level Backup Restore
(days) tapes
1 0  n/a  1
2 3  1  1, 2
3 2  2  1, 3
4 5  1  1, 2, 4
5 4  2  1, 2, 5
6 7  1  1, 2, 5, 6
7 6  2  1, 2, 5, 7
8 9  1  1, 2, 5, 7, 8
9 8  2  1, 2, 5, 7, 9
10 9  1  1, 2, 5, 7, 9, 10
11 9  1  1, 2, 5, 7, 9, 10, 11
… 9  1  1, 2, 5, 7, 9, 10, 11, …

10.6 What to back up

You want to back up as much as possible. The major exception is software that can be easily reinstalled,^{4} but even they may have configuration files that it is important to back up, lest you need to do all the work to configure them all over again. Another major exception is the /proc filesystem; since that only contains data that the kernel always generates automatically, it is never a good idea to back it up. Expecially the /proc/kcore file is unnecessary, since it is just an image of your current physical memory; it’s pretty large as well.

Gray areas include the news spool, log files, and many other things in /var. You must decide what you consider important.

The obvious things to back up are user files (/home) and system configuration files (/etc, but possibly other things scattered all over the filesystem).

10.7 Compressed backups

Backups take a lot of space, which can cost quite a lot of money. To reduce the space needed, the backups can be compressed. There are several ways of doing this. Some programs have support for for compression built in; for example, the –gzip (-z) option for GNU tar pipes the whole backup through the gzip compression program, before writing it to the backup medium.

4You get to decide what’s easy. Some people consider installing from dozens of floppies easy.

Unfortunately, compressed backups can cause trouble. Due to the nature of how compression works, if a single bit is wrong, all the rest of the compressed data will be unusable. Some backup programs have some built in error correction, but no method can handle a large number of errors. This means that if the backup is compressed the way GNU tar does it, with the whole output compressed as a unit, a single error makes all the rest of the backup lost. Backups must be reliable, and this method of compression is not a good idea.

An alternative way is to compress each file separately. This still means that the one file is lost, but all other files are unharmed. The lost file would have been corrupted anyway, so this situation is not much worse than not using compression at all. The afio program (a variant of cpio) can do this.

Compression takes some time, which may make the backup program unable to write data fast enough for a tape drive.^{5} This can be avoided by buffering the output (either internally, if the backup program if smart enough, or by using another program), but even that might not work well enough. This should only be a problem on slow computers.

5If a tape drive doesn’t data fast enough, it has to stop; this makes backups even slower, and can be bad for the tape and the drive.

Keeping Time

Time is an illusion. Lunchtime double so. (Douglas Adams.)

This chapter explains how a Linux system keeps time, and what you need to do to avoid causing trouble. Usually, you don’t need to do anything about time, but it is good to understand it.

11.1 Time zones

Time measurement is based on mostly regular natural phenomena, such as alternating light and dark periods caused by the rotation of the planet. The total time taken by two successive periods is constant, but the lengths of the light and dark period vary. One simple constant is noon.

Noon is the time of the day when the Sun is at its highest position. Since the Earth is round,^{1} noon happens at different times in different places. This leads to the concept of local time. Humans measure time in many units, most of which are tied to natural phenomena like noon. As long as you stay in the same place, it doesn’t matter that local times differ.

As soon as you need to communicate with distant places, you’ll notice the need for a common time. In modern times, most of the places in the world communicate with most other places in the world, so a global standard for measuring time has been defined. This time is called universal time (UT or UTC, formerly known as Greenwich Mean Time or GMT, since it used to be local time in Greenwich, England). When people with different local times need to communicate, they can express times in universal time, so that there is no confusion about when things should happen.

1According to recent research.

Each local time is called a time zone. While geography would allow all places that have noon at the same time have the same time zone, politics makes it difficult. For various reasons, many countries use daylight savings time, that is, they move their clocks to have more natural light while they work, and then move the clocks back during winter. Other countries do not do this. Those that do, do not agree when the clocks should be moved, and they change the rules from year to year. This makes time zone conversions definitely non-trivial.

Time zones are best named by the location or by telling the difference between local and universal time. In the US and some other countries, the local time zones have a name and a three letter abbreviation. The abbreviations are not unique, however, and should not be used unless the country is also named. It is better to talk about the local time in, say, Helsinki, than about East European time, since not all countries in Eastern Europe follow the same rules.

Linux has a time zone package that knows about all existing time zones, and that can easily be updated when the rules change. All the system administrator needs to do is to select the appropriate time zone. Also, each user can set his own time zone—this is important since many people work with computers in different countries over the Internet. When the rules for daylight savings time change in your local time zone, make sure you’ll upgrade at least that part of your Linux system. Other than setting the system time zone and upgrading the time zone data files, there is little need to bother about time.

11.2 The hardware and software clocks

A personal computer has a battery driven hardware clock. The battery ensures that the clock will work even if the rest of the computer is without electricity. The hardware clock can be set from the BIOS setup screen or from whatever operating system is running.

The Linux kernel keeps track of time independently from the hardware clock. During the boot, Linux sets its own clock to the same time as the hardware clock. After this, both clocks run independently. Linux maintains its own clock because looking at the hardware is slow and complicated.

The kernel clock always shows universal time. This way, the kernel does not need to know about time zones at all—the simplicity results in higher reliability and makes it easier to update the time zone information. Each process handles time zone conversions itself (using standard tools that are part of the time zone package).

The hardware clock can be in local time or in universal time. It is usually better to have it in universal time, because then you don’t need to change the hardware clock when daylight savings time begins or ends (UTC does not have DST). Unfortunately, some PC operating systems—including MS-DOS, Windows, OS/2—assume the hardware clock shows local time. Linux can handle either, but if the hardware clock shows local time, then it must be modified when daylight savings time begins or ends (otherwise it wouldn’t show local time).

11.3 Showing and setting time

In the Debian system, the system time zone is determined by the symbolic link /etc/localtime. This link points at a time zone data file that describes the local time zone. The time zone data files are stored in /usr/lib/zoneinfo. Other Linux distributions may do this differently.

A user can change his private time zone by setting the TZ environment variable. If it is unset, the system time zone is assumed. The syntax of the TZ variable is described in the tzset(3) manual page.

The date command shows the current date and time.^{2} For example:

     $ date
     Sun Jul 14 21:53:41 EET DST 1996
     $

That time is Sunday, 14th of July, 1996, at about ten before ten at the evening, in the time zone called “EET DST” (which might be East European Daylight Savings Time). date can also show the univeral time:

     $ date -u
     Sun Jul 14 18:53:42 UTC 1996
     $

date is also used to set the kernel’s software clock:

2Beware of the time command, which does not show the current time.

     # date 07142157
     Sun Jul 14 21:57:00 EET DST 1996
     # date
     Sun Jul 14 21:57:02 EET DST 1996
     #

See the date manual page for more details—the syntax is a bit arcane. Only root can set the time. While each user can have his own time zone, the clock is the same for everyone.

date only shows or sets the software clock. The clock commands syncronizes the hardware and software clocks. It is used when the system boots, to read the hardware clock and set the software clock. If you need to set both clocks, you first set the software clock with date, and then the hardware clock with clock -w.

The -u option to clock tells it that the hardware clock is in universal time. You must use the -u option correctly. If you don’t, your computer will be quite confused about what the time is.

The clocks should be changed with care. Many parts of a Unix system require the clocks to work correctly. For example, the cron daemon runs commands periodically. If you change the clock, it can be confused of whether it needs to run the commands or not. On one early Unix system, someone set the clock twenty years into the future, and cron wanted to run all the periodic commands for twenty years all at once. Current versions of cron can handle this correctly, but you should still be careful. Big jumps or backward jumps are more dangeours than smaller or forward ones.

11.4 When the clock is wrong

The Linux software clock is not always accurate. It is kept running by a periodic timer interrupt generated by PC hardware. If the system has too many processes running, it may take too long to service the timer interrupt, and the software clock starts slipping behind. The hardware clock runs independently and is usually more accurate. If you boot your computer often (as is the case for most systems that aren’t servers), it will usually keep fairly accurate time.

If you need to adjust the hardware clock, it is usually simplest to reboot, go into the BIOS setup screen, and do it from there. This avoids all trouble that changing system time might cause. If doing it via BIOS is not an option, set the new time with date and clock (in that order), but be prepared to reboot, if some part of the system starts acting funny.

A networked computer (even if just over the modem) can check its own clock automatically, by comparing it to some other computer’s time. If the other computer is known to keep very accurate time, then both computers will keep accurate time. This can be done by using the rdate and netdate commands. Both check the time of a remote computer (netdate can handle several remote computers), and set the local computer’s time to that. By running one these commands regularly, your computer will keep as accurate time as the remote computer.

META: say something intelligent about NTP

Appendix A

Measuring Holes

This appendix contains the interesting part of the program used to measure the potential for holes in a filesystem. The source distribution of the book contains the

full source code (sag/measure-holes/measure-holes.c).
int process(FILE *f, char *filename) {
static char *buf = NULL;
static long prev_block_size = -1;
long zeroes;
char *p;
if (buf == NULL || prev_block_size != block_size) {
        free(buf);
        buf = xmalloc(block_size + 1);
        buf[block_size] = 1;
        prev_block_size = block_size;
}
zeroes = 0;
while (fread(buf, block_size, 1, f) == 1) {
        for (p = buf; *p == '\0'; )
                ++p;
        if (p == buf+block_size)
                zeroes += block_size;
}
if (zeroes > 0)
        printf("%ld %s\n", zeroes, filename);
if (ferror(f)) {
        errormsg(0, -1, "read failed for `%s'", filename);
        return -1;
}
return 0;
}

Appendix B

Glossary (DRAFT)

The Librarian of the Unseen University had unilaterally decided to aid comprehension by producing an Orang-utan/Human Dictionary.

He’d been working on it for three months.

It wasn’t easy. He’d got as far as ‘Oook.’

(Terry Pratchett, “Men At Arms”)

This is a short list of word definitions for concepts relating to Linux and system administration. The page references are to the first or most important place where the word is used.

ambition The act of writing funny sentences in the hope of getting them into the Linux cookie file.

application program (p. 6) Software that does something useful. The results of using an application program is what the computer was bought for. See also system program, operating system.

daemon A process lurking in the background, usually unnoticed, until something triggers it into action. For example, the update daemon wakes up every thirty seconds or so to flush the buffer cache, and the sendmail daemon awakes whenever someone sends mail.

file system (p. 38) The methods and data structures that an operating system uses to keep track of files on a disk or partition; the way the files are organized on the disk. Also used about a partition or disk that is used to store the files or the type of the filesystem.

glossary A list of words and explanations of what they do. Not to be confused with a dictionary, which is also a list of words and explanations.

kernel (p. 6) Part of an operating system that implements the interaction with hardware and the sharing of resources. See also system program.

operating system (p. 6) Software that shares a computer system’s resources (processor, memory, disk space, network bandwidth, and so on) between users and the application programs they run. Controls access to the system to provide security. See also kernel, system program, application program.

system call (p. 6) The services provided by the kernel to application programs, and the way in which they are invoked. See section 2 of the manual pages.

system program (p. 6) Programs that implement high level functionality of an operating system, i.e., things that aren’t directly dependent on the hardware. May sometimes require special privileges to run (e.g., for delivering electronic mail), but often just commonly thought of as part of the system (e.g., a compiler). See also application program, kernel, operating system.

Bibliography

[Anv] Peter Anvin. Linux device list. A list of major and minor device numbers for Linux devices. Now included in the kernel sources.

[Cha] Graham Chapman. Bootdisk howto. Available with other Linux HOWTO’s.

[Kir] Olaf Kirch. Linux network administrators’ guide.

[Qui95] Daniel Quinlan. Linux Filesystem Structure—Release 1.2, March 1995. A description of and a proposal for a standard Linux directory tree, with the intention is to make it easier to package software and administer Linux systems by making files appear in standard places. Follows fairly closely traditional Unix practice, and has got support from most Linux distributions. Available

via FTP from ftp.funet.fi, directory /pub/Linux/doc/fsstnd.

[TV] Stephen Tweedie and Alexei Vovenko. Linux filesys-

tem  defragmenter.  Available  electronically  from

ftp://sunsite.unc.edu/pub/Linux/system/Filesystems/defrag-0.6.tar.gz.

[Wel] Matt Welsh. Installation and getting started guide.

Index

.hushlogin, 80 .profile, 82, 86 /bin, 12, 17, 20, 53 /bin/sh, 46, 72, 82 /boot, 17 /dev, 12, 17, 20, 25, 53 /dev/MAKEDEV, 20 /dev/MAKEDEV.local, 20 /dev/fd0, 29, 32 /dev/fd0H1440, 29 /dev/fd1, 29 /dev/hda, 28 /dev/hda1, 38 /dev/hda2, 44 /dev/hdb, 28 /dev/hdc, 28 /dev/hdd, 28 /dev/sda, 26, 28 /dev/sdb, 28 /dev/sdb7, 38 /dev/tty, 73 /dev/tty1, 74 /etc, 12, 17, 18, 53, 98 /etc/csh.cshrc, 19 /etc/csh.login, 19 /etc/fastboot, 48 /etc/fdprm, 18, 29 /etc/fstab, 18, 46, 47, 55, 59, 60 /etc/group, 18, 81, 87, 88 /etc/init.d/rc, 75 /etc/inittab, 18, 73–75 /etc/issue, 18, 78 /etc/localtime, 102 /etc/login.defs, 19 /etc/magic, 18 /etc/motd, 19, 80 /etc/mtab, 19 /etc/nologin, 80 /etc/passwd, 18, 19, 77, 81, 85–88 /etc/printcap, 19 /etc/profile, 19, 82, 87 /etc/rc, 18 /etc/rc.d, 18 /etc/rc?.d, 18 /etc/securetty, 19, 81 /etc/shadow, 19, 81, 86 /etc/shadow.group, 81 /etc/shells, 19 /etc/skel, 86, 87 /etc/skel/.profile, 86 /etc/termcap, 19 /extra-swap, 58 /home, 12, 15, 16, 18, 44, 53–55, 98 /home/staff, 16 /home/students, 16 /lib, 12, 17, 53 /lib/modules, 17 /mnt, 17 /mnt/dosa, 17 /mnt/exta, 18 /proc, 18, 22, 23, 40, 98 /proc/1, 22 /proc/cpuinfo, 22 /proc/devices, 22 /proc/dma, 22 /proc/filesystems, 22 /proc/interrupts, 22 /proc/ioports, 22 /proc/kcore, 22, 41, 98 /proc/kmsg, 23 /proc/ksyms, 23 /proc/loadavg, 23 /proc/meminfo, 23, 59 /proc/modules, 23 /proc/net, 23 /proc/self, 23 /proc/stat, 23 /proc/uptime, 23 /proc/version, 23 /root, 17 /sbin, 17 /sbin/getty, 9 /sbin/init, 9, 67, 72 /sbin/update, 63 /tmp, 10, 17, 22, 45, 53, 73, 76 /usr, 12, 15, 16, 18, 20, 44, 45, 53–55,

76 /usr/X11R6, 20 /usr/X386, 20 /usr/adm/messages, 33 /usr/bin, 20 /usr/doc, 20 /usr/include, 20 /usr/info, 20 /usr/lib, 20, 21 /usr/lib/libc.a, 16 /usr/lib/zoneinfo, 102 /usr/local, 20, 21 /usr/local/bin, 20 /usr/man, 20 /usr/man/cat*, 21 /usr/man/man*, 21 /usr/sbin, 20 /usr/var, 16 /var, 12, 15, 16, 18, 21, 45, 98 /var/adm/messages, 16 /var/catman, 21 /var/lib, 21 /var/local, 21 /var/lock, 21 /var/log, 21 /var/log/messages, 21 /var/log/wtmp, 21, 80 /var/man, 21 /var/run, 21 /var/run/utmp, 21, 80 /var/spool, 22 /var/spool/mail, 12, 22 /var/tmp, 10, 17, 22, 45 /vmlinuz, 17

active partition, 66

adduser, 85 afio, 99 application program, 6 at, 10, 88 Athena, 10 axe, 69

background processes

and shutting down, 68 backups, 70 badblocks, 33, 43, 49 bdflush, 50, 63 BIOS, 65 boot process, 9 boot record

master, see MBR

partition, see partition boot record boot sector, 65 booting, 64

from emergency floppy, 70 from floppy, 66 from hard disk, 66 messages, 67

bootstrap loader, 65 BSD, 8 buffer cache, 7, 65, 68, 69

C compiler, 7 cache, 68 centralized computing, 11 cfdisk, 36 checking a filesystem, 67 chfn, 88 chmod, 87, 88 chown, 87 chsh, 19, 88 cleaning temporary files, 10 clock, 103 communication, 12 compression of kernel, 67 console, 68 cpio, 92, 93, 97, 99 crack, 81 crash, 69 cron, 10, 80, 88, 103 crontab, 10 crying, 64 ctrl-alt-del, 70 curs_termcap, 19

daemons at, see at cron, see cron init, see /sbin/init rlogin, see rlogin starting, 68

telnet, see telnet data loss, 65 date, 102, 103 dd, 52 debugfs, 48, 51 deluser, 89 device driver, 7 device driver configuration, 67 df, 19, 50 disk cache, 68 distributed computing, 11 documentation, 7 DouBle, 55 du, 50 dump, 47, 51, 92, 93, 97 dumpe2fs, 51

e-mail, see electronic mail e2fsck, 48–50 electronic mail, 12 elm, 12 emergency floppy

making, 71 using, 70

emergency shutdown, see shutdown, emer-

gency error messages, 10 exec, 78

fault tolerance, 11 fdformat, 32, 33 fdisk, 34, 36, 37 fdisk -l, 34 file, 18 filesystem driver, 7 Filesystem Standard, see FSSTND filesystems

root, 67 find, 88, 95 fips, 38 floppy

booting from, see booting floppy drive, 70 floppy-image, 52 fork, 78 free, 23, 59–61 fsck, 33, 45, 47–49, 51, 67, 76, 77 FSSTND, 13 fstab, 46 ftpd, 19

games, 7 GCC, 7 getty, 9, 11, 18, 68, 72–74, 78, 79, 83 graphical user interface, 10 group, 18 gzexe, 55 gzip, 55, 99

hard disk

booting from, see booting hardware configuration, 67

improper shutdown, see shutdown, emer-

gency incoming mailbox, 12 inetd, 80 init, 9, 18, 65, 67–69, 72, see /sbin/init,

73–76, 78, 79, 83 and ctrl-alt-del, 70 single user mode, 70

inittab, 74, 76 installation disks, 70

kernel, 6 kernel compression, 67 kernel configuration, 67 kernel panic, 67 kill -HUP 1, 73 killing processes, 9

last, 80 letter, see electronic mail LILO, 66, 67 Linux Filesystem Standard, see FSSTND log files, 10 logging in, 9 logging out, 9 login, 9, 19, 21, 78, 80, 83 lpr, 25 ls, 25, 26, 62 ls -l, 25 lseek, 39

magic, 18 mail, electronic, see electronic mail mailbox, 12 MAKEDEV, 20 master boot record, 65 MBR, 65 memory management, 7 mkdir, 87 mke2fs, 50 mkfs, 33, 42, 43 mkfs), 2 mkswap, 57, 58 Motif, 10 mount, 18, 19, 44, 45, 47 mount -a, 47 mounting

root filesystem, 67 mounting filesystems, 9 mtools, 46 multitasking, 7 multiuser mode, 9

netdate, 104 nethack, 76 Network File System, see NFS network file systems, 11 network management, 7 NFS, 11 now, 50

Open Look, 10 operating system, 6

panic, 67 partition boot record, 65 partition table, 66 passwd, 18, 81, 85, 87, 88 periodic command execution, see cron

and at pine, 12 printer queue, 12 proc, 22 process management, 7 programming languages, 7 ps, 61

ramdisk, 71 rdate, 104 rdev, 45, 67 read-only root filesystem, 67 rebooting, 70 restore, 51 rlogin, 11, 79 root, 9, 46 root filesystem, 67 run level, 9

sendmail, 12, 106 serial terminal, 68 setfdprm, 18, 29, 32 sharing printers, 12 shutdown, 68–70, 80

actions, 69 emergency, 69 triple syncs, 69 using, 68

shutdown message, 69 shutting down, 68 single user mode, 9, 70 smail, 12 social relations, 12 socket, 8 spooling, 12 starting a computer, 64 startup messages, 67 stopping the processor, 9 su, 81, 89 sudo, 46 swap space, 7 swapoff, 59 swapon, 58, 59 sync, 50, 62, 63, 69 syslog, 9, 21, 23, 33, 80 system call, 6 system crash, 69 system program, 6 sysvinit, 73

tail, 89 tar, 52, 92–97, 99 telinit, 75, 76 telnet, 7, 11, 79 temporary files, 10 termcap, 19 terminal, 68 terminfo, 19 text mode, 67 time, 102 top, 59 tune2fs, 50, 51 tzset, 102

umask, 47 umount, 46, 47 unmount, 46 unmounting filesystems, 9 update, 50, 63, 69, 106 user authentication, 9 user mode, 7 useradd, 85 userdel, 89 utmp, 80

video card, 67 video mode, 67 vigr, 87, 88 vipw, 87, 88 virtual console, 68 virtual filesystem, 8

w, 80 warning messages, 10 who, 80 Windows 95, see crying wtmp, 80

X Window System, 9, 10 xterm, 69

zip, 55